### v1

# Asymptotic Stability of 3D Out-flowing Compressible Viscous Fluid under Non-Spherical Perturbation

Abstract
We study an outflow problem for the $3$-dimensional isentropic compressible Navier-Stokes equations. The fluid under consideration occupies the exterior domain of unit ball, $\Omega=\{x\in\mathbb{R}^{3}\,|\,|x|\geq 1\}$, and it is flowing out from $\Omega$ at a constant speed $|u_{b}|$, in the normal direction to the boundary surface $\partial\Omega$. It is shown in Hashimoto-Matsumura(2021) that if the fluid velocity at the far-field is assumed to be zero, and $|u_{b}|$ is sufficiently small, then there exists a unique spherically symmetric stationary solution $(\tilde{\rho},\tilde{u})(x)$. In this paper, we prove that $(\tilde{\rho},\tilde{u})(x)$ is asymptotically stable in time, under small, possibly non-spherically symmetric initial perturbations.
Keywords.
Navier–Stokes equation, Outflow Problem, Stationary wave,
AMS subject classifications.
35B35, 35B40, 76N15.
1 Introduction
The Navier-Stokes equations for the isentropic motion of compressible viscous gas in the Eulerian coordinate is the system of equations given by
$\partial_{t}\rho+\textnormal{{div}}(\rho u)=0,$ (1.1a)
$\rho\{\partial_{t}u+(u\cdot\nabla)u\}+\nabla P(\rho)=\mu\Delta u+(\mu+\lambda)\nabla\textnormal{{div}}u\equiv Lu.$ (1.1b)
We study the asymptotic behaviour of a solution $(\rho,u)$ to (1.1) in an unbounded exterior domain $\Omega\vcentcolon=\{x=(x_{1},x_{2},x_{3})^{\top}\in\mathbb{R}^{3}\;|\;|x|>1\}$. Here $\rho>0$ is the mass density; $u=(u^{1},u^{2},u^{3})^{\top}$ is the velocity of gas; $P(\rho)=K\rho^{\gamma}\;(K>0,\gamma\geq 1)$ is the pressure with the adiabatic exponent $\gamma$; $\mu$ and $\lambda$ are constant called viscosity-coefficients satisfying $\mu>0$ and $2\mu+3\lambda\geq 0$.
In this paper, we consider the model where fluid is flowing out at a constant velocity from a solid inner sphere centred at the origin of radius $|x|=1$. Its corresponding boundary condition is given as:
[…]
For equations (1.1)–(1.2), we prescribe the initial data $(\rho_{0},u_{0})(x)$ such that
$\rho(x,0)=\rho_{0}(x)>0,\quad u(x,0)=u_{0}(x),$ (1.3a)
$\lim_{|x|\to\infty}(\rho_{0}(x),u_{0}(x))=(\rho_{+},0),\quad 0<\rho_{+}<\infty.$ (1.3b)
In addition, (1.3a) is compatible with the boundary data (1.2a) which states that,
$u_{0}|_{x\in\partial\Omega}=u_{b}\frac{x}{|x|}\big{|}_{x\in\partial\Omega},$ (1.4a)
$\big{\{}-\rho_{0}(u_{0}\cdot\nabla)u_{0}+\mu\Delta u_{0}+(\mu+\lambda)\nabla\textnormal{{div}}u_{0}-\nabla P(\rho_{0})\big{\}}\big{|}_{x\in\partial\Omega}=0.$ (1.4b)
Since the characteristic speed of (1.1a) is negative on the boundary due to (1.2a), no boundary condition for $\rho$ at $x\in\partial\Omega$ is necessary for the well-posedness of the initial boundary value problem (1.1)–(1.4).
Spehrically symmetric steady-state solution
Consider instead the general multi-dimensional exterior domain $\Omega=\{x\in\mathbb{R}^{n}\,|\,|x|\geq 1\}$ where $n\geq 2$. Let $(\tilde{\rho},\tilde{u})$ be the steady state solution of (1.1)–(1.4), solving the system:
$\textnormal{{div}}\big{(}\tilde{\rho}\tilde{u}\big{)}=0$ for $\text{for }\ x\in\Omega,$ (1.5a)
$\tilde{\rho}(\tilde{u}\cdot\nabla)\tilde{u}+\nabla P(\tilde{\rho})=L\tilde{u}$ for $\text{for }\ x\in\Omega,$ (1.5b)
$\tilde{u}(x)=u_{b}\frac{x}{|x|}$ for $\text{for }\ x\in\partial\Omega,$ (1.5c)
$(\tilde{\rho},\tilde{u})(x)\to(\rho_{+},0)$ as |$\text{as }\ |x|\to\infty.$ (1.5d)
Since the boundary and far-field conditions (1.5c)–(1.5d) are spherically symmetric, if the unique solution $(\tilde{\rho},\tilde{u})$ to (1.5) exists, then $(\tilde{\rho},\tilde{u})$ must also be spherically symmetric. This means that $(\tilde{\rho},\tilde{u})$ can be expressed in the form:
$\tilde{\rho}(x)=\tilde{\rho}(r),\quad\tilde{u}(x)=\tilde{U}(r)\frac{x}{r},\qquad\text{where }\ r\equiv|x|\geq 1,$ (1.6)
for some function $\tilde{U}(r)\vcentcolon[1,\infty)\to\mathbb{R}$. Then the system (1.5) is reformulated as
$\partial_{r}\big{(}r^{n-1}\tilde{\rho}\tilde{U}\big{)}=0$ for $\text{for }\ r\geq 1,$ (1.7a)
$\tilde{\rho}\tilde{U}\partial_{r}\tilde{U}+\partial_{r}P(\tilde{\rho})=(2\mu+\lambda)\partial_{r}\Big{(}\frac{(r^{n-1}\tilde{U})_{r}}{r^{n-1}}\Big{)}$ for $\text{for }\ r\geq 1,$ (1.7b)
$\tilde{U}(1)=u_{b},\quad\lim_{r\to\infty}\big{(}\tilde{\rho}(r),\tilde{U}(r)\big{)}=(\rho_{+},0).$ (1.7c)
Multiplying $r^{n-1}$ on (1.5a) and integrating the resultant equality over $(1,r)$, we obtain
$\tilde{U}(r)=\dfrac{u_{b}\tilde{\rho}(1)}{r^{n-1}\tilde{\rho}(r)},$
for $r\geq 1$. If a solution $(\tilde{\rho},\tilde{U})(r)$ exists and $\tilde{\rho}(r)\to\rho_{+}\in(0,\infty)$ as $r\to\infty$, then it is necessary to have the far-field condition, $\tilde{U}(r)\to 0$ as $r\to\infty$.
For general spatial dimension $n\geq 2$, I. Hashimoto and A. Matsumura in [3] proved the existence and uniqueness of a spherically symmetric classical solution to (1.5), provided that $|u_{b}|$ is small. Subsequently in [4], I. Hashimoto, S. Nishibata and S. Sugizaki refined the spatial decay estimate for $(\tilde{\rho},\tilde{U})$, and they also prove the asymptotic stability of $(\tilde{\rho},\tilde{U})$ for small, spherically symmetric initial perturbations. In particular, the well-posedness theorem for the stationary solution can be summarised as follows:
Lemma 1.1.
There exists $\delta=\delta(\rho_{+},\mu,\gamma,K,n)>0$ such that if $|u_{b}|\leq\delta$, then the unique solution $(\tilde{\rho},\tilde{U})(r)\in\mathcal{C}^{2}[1,\infty)$ to the problem (1.7) exists. In addition, $\tilde{\rho}(r)$ and $\tilde{U}(r)$ are respectively strictly monotone increasing and decreasing. Moreover, there exists constant $C=C(\rho_{+},\mu,\gamma,K,n)>0$ such that for $r\geq 1$,
$\tilde{U}(r)=u_{b}\tilde{\rho}(1)\dfrac{r^{1-n}}{\tilde{\rho}(r)},\quad$ $|\tilde{\rho}(r)-\rho_{+}|\leq Cr^{-2n+2},$
$C^{-1}|u_{b}|r^{-n}\leq|\partial_{r}\tilde{U}(r)|\leq C|u_{b}|r^{-n},\quad$ $C^{-1}|u_{b}|^{2}r^{-2n+1}\leq|\partial_{r}\tilde{\rho}(r)|\leq C|u_{b}|^{2}r^{-2n+1},$
$C^{-1}|u_{b}|r^{-2n}\leq|\partial_{r}^{2}\tilde{U}(r)|\leq C|u_{b}|r^{-2n},\quad$ $C^{-1}|u_{b}|^{2}r^{-2n}\leq|\partial_{r}^{2}\tilde{\rho}(r)|\leq C|u_{b}|^{2}r^{-2n}.$
Y. Huang and S. Nishibata in [6] prove that if the initial data (1.3) is spherically symmetric, then the stationary solution $(\tilde{\rho},\tilde{u})$ obtained in Lemma 1.1 is still asymptotically stable even if the initial perturbation is large. However, the stability under general possibly non-spherically symmetric initial perturbation remained an open question. In this paper, we give a positive answer to this problem provided that the initial perturbation $(\rho_{0}-\tilde{\rho},u_{0}-\tilde{u})$ is small in some suitable Sobolev space.
2 Main Theorem
Theorem 2.1.
There exists $\varepsilon=\varepsilon(\rho_{+},\mu,\gamma,K,n)>0$ such that if
$|u_{b}|+\|(\rho_{0}-\tilde{\rho},u_{0}-\tilde{u})\|_{H^{3}(\Omega)}\leq\varepsilon,$
then the initial boundary value problem (1.1)–(1.4) has a unique solution $(\rho,u)$ globally-in-time, which satisfies
$\rho-\tilde{\rho}\in\mathcal{C}^{0}\big{(}[0,\infty);H^{3}(\Omega)\big{)}\cap\mathcal{C}^{1}\big{(}[0,\infty);H^{2}(\Omega)\big{)}\cap L^{2}\big{(}0,\infty;H^{3}(\Omega)\big{)},$
$u-\tilde{u}\in\mathcal{C}^{0}\big{(}[0,\infty);H^{3}(\Omega)\big{)}\cap\mathcal{C}^{1}\big{(}[0,\infty);H^{1}(\Omega)\big{)}\cap L^{2}\big{(}0,\infty;H^{4}(\Omega)\big{)}.$
In addition, the following convergence holds,
$\lim\limits_{t\to\infty}\sup\limits_{x\in\Omega}|(\rho-\tilde{\rho},u-\tilde{u})(t,x)|=0.$
Remark 2.2.
The out-flowing boundary condition (1.2) plays a significant role in the derivation for $L^{\infty}_{T}H^{2}$-estimates of $\big{(}\nabla(\rho-\tilde{\rho}),\nabla(u-\tilde{u})\big{)}$. Since $\partial\Omega$ is a sphere, we observe that the temporal and angular derivatives preserve the boundary condition (1.2), while the radial derivative of $u$ at $\partial\Omega$ is not a-priori given. Due to this, it is necessary to split the analysis into 3 steps as follows: 1. obtain estimates for temporal and angular derivatives of $(\rho-\tilde{\rho},u-\tilde{u})$ using the parabolic structure of (1.1b); 2. obtain estimates for radial derivative of $\rho-\tilde{\rho}$ using the damping effect of compressible Navier-Stokes equations; 3. obtain estimates for full spatial derivatives of $(\rho-\tilde{\rho},u-\tilde{u})$ by treating the system (1.1) as Stokes equations. It is worth noting that in step 2 (Section 3.4), a cancellation occurs on the radial derivatives of $u$. This serves as the key argument to successfully close the estimate.
Related results.
The compressible Navier-Stokes equation is an important subject in mathematical physics. Particularly, the problems of out-flowing or in-flowing boundary have gained traction in recent decades. We state several previous results, which are relevant to the present paper.
First, for a comprehensive survey of the mathematical theory of compressible Navier-Stokes equation, we refer readers to the book [7] by S. N. Antontsev, A. V. Kazhikhov, and V. N. Monakhov. The first breakthrough research on the large time stability of solution is done by A. Matsumura and T. Nishida in [15], where they consider the heat-conductive compressible flow for a general $3$-dimensional exterior domain with adhesion boundary condition ($u|_{\partial\Omega}=0$ and $\partial_{n}\theta|_{\partial\Omega}=0$ where $\theta$ is the absolute temperature). In this case the stationary solution is $(\tilde{\rho},0,\bar{\theta})$ where $\tilde{\rho}=\tilde{\rho}(x)$ is a positive function of spatial variable, and $\bar{\theta}$ is a positive constant. Under smallness assumptions on the initial perturbation and external forces, it is proved that $(\tilde{\rho},0,\bar{\theta})$ is stable globally in time.
When the equation is spherically symmetric with adhesion boundary condition, a pioneering work has been done by N. Itaya [9], which establishes the global-in-time existence of a unique classical solution on a bounded annulus domain. In this work, no smallness assumption on the initial data is imposed. Later, T. Nagasawa studies the asymptotic state for the same problem in [20]. The paper [9] has motivated a sequence of developments on the topic of spherically symmetric solution. For instance, A. Matsumura in [12] constructed a spherically symmetric classical solution to the isothermal model with external forces on a bounded annular domain. Moreover, he also shows the convergence rate to the stationary solution as time tends to infinity is exponential. Subsequently, this result has been extended to the isentropic and heat-conductive models by K. Higuchi in [8]. The well-posedness of spherically symmetric solution in an unbounded exterior domain is first obtained by S. Jiang [10], where the global-in-time existence of a uniqueness classical solution is shown. In addition, a partial result on the time asymptotic stability is proved in [10] where, for $n=3$, $\|u(\cdot,t)\|_{L^{2j}}\to 0$ as $t\to\infty$ with any fixed integer $j\geq 2$. Later, this restriction on the long time stability was fully resolved by T. Nakamura and S. Nishibata in [18], where a complete stability theorem was obtained for both $n=2$ and $3$ with large initial data. We also refer to the paper [17] by T. Nakamura, S. Nishibata, and S. Yanagi, where the time asymptotic stability of the spherically symmetric solution for the isentropic flow is established with large initial data and external forces.
For the general outflow or inflow problem, the stationary solution becomes non-trivial. This leads to a variety of physically interesting time-asymptotic behaviours for the solutions. A. Matsumura in [13] starts the first investigation of this problem for the isentropic model posed on the one dimensional half-space domain. Several kinds of boundary conditions were considered in [13], which includes inflow, outflow, and adhesion boundary conditions. He formulated conjectures on the classification of asymptotic behaviours of the solutions in different cases subject to the relation between the boundary data and the spatial asymptotic data. Then the stability theorems for some cases of inflow problem were established by A. Matsumura and K. Nishihara in [16] where they employed the Lagrangian mass coordinate. Following this work, S. Kawashima, S. Nishibata and P. Zhu in [11] further examine the outflow problem in one dimensional half space. They prove the long time stability of solutions with small initial perturbation with respect to the stationary solution. A detailed desicription of the convergence rate towards the stationary solution is found in the work of T. Nakamura, S. Nishibata and T. Yuge [17].
For the non-isentropic inflow problem in the half line, T. Nakamura and S. Nishibata in [19] demonstrate the time-asymptotic stability of stationary solutions under a small initial perturbation, for both the subsonic and transonic cases.
The research on the outflow and inflow problems for the spherically symmetric solution in an unbounded exterior domain has been relatively new subject. Under the assumption that velocity at boundary, $|u_{b}|$ is sufficiently small, I. Hashimoto and A. Matsumura in [3] employ the iteration method to obtain the existence of a unique spherically symmetric stationary solutions for both inflow and outflow problems in an exterior domain. More recently, I. Hashimoto, S. Nishibata, S. Sugizaki in [4] show the stability of spherically symmetric stationary solution for outflow problem under small initial perturbation. Subsequently, Y. Huang and S. Nishibata in [6] considers the large, spherically symmetric initial perturbation for the outflow problem and they prove that the stationary solution remains stable in large time without any smallness assumption on the initial data. The main theorem proved in this paper is the first stability result for outflow problem under general possibly non-spherically symmetric initial perturbation.
Outline of the paper.
At the beginning of Section 3, we restate the equations (1.1) in terms of the difference functions $(\phi,\psi)=(\rho-\tilde{\rho},u-\tilde{u})$, and we reformulate them in linear forms which is suitable for the a-priori estimates. Moreover, we also provide the precise definition of energy norm $N(T)$ that is used throughout the proof. Then Section 3 is divided into 5 subsections as follows: 1. in Section 3.1, we derive the $L^{2}$–estimates of $(\phi,\psi)$ through the use of $L^{2}$ relative energy density, denoted $\mathcal{E}$; 2. in Sections 3.23.3, the estimates for temporal and angular derivatives are obtained using the parabolic structure of the momentum equation (1.1b); 3. in Section 3.4, we derive estimates involving the radial derivative of $\phi$, which is achieved via the coupling structure of continuity equation (1.1a) and momentum equation (1.1b). Finally in Section 3.5, we obtain estimates for the full derivatives of $(\phi,\psi)$, by treating (1.1b) as an elliptic system, and (1.1a)–(1.1b) as a Stokes-type system.
Notations.
  1. 1.
    For $T\geq 0$, we denote the space-time domains as
    $Q_{T}\vcentcolon=[0,T]\times\Omega,\qquad\Gamma_{T}\vcentcolon=[0,T]\times\partial\Omega.$
  2. 2.
    For Cartesian coordinate $(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$, we denote $\hat{e}_{1}$, $\hat{e}_{2}$, $\hat{e}_{3}$ as the corresponding unit basis vectors. Moreover, for a scalar valued function $F\vcentcolon\mathbb{R}^{3}\to\mathbb{R}$ and a vector-valued function $V\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$ with $V=(V^{1},V^{2},V^{3})^{\top}$, we denote the differential operators:
    $\nabla F\vcentcolon=\sum_{i=1}^{3}\hat{e}_{i}\partial_{x_{i}}F,\quad$ $\textnormal{{div}}V\vcentcolon=\sum_{i=1}^{3}\partial_{x_{i}}V^{i},\quad$ $\Delta F\vcentcolon=\sum_{i=1}^{3}\partial_{x_{i}}^{2}F,$
    $(V\cdot\nabla)F\vcentcolon=\sum_{i=1}^{3}V^{i}\partial_{x_{i}}F,\quad$ $D_{t}F\vcentcolon=\partial_{t}F+(u\cdot\nabla)F.$
    In addition, for operators $S$, $T$ we denote $[S,T]\vcentcolon=ST-TS$ as the commutator.
  3. 3.
    For $k>0$, the Sobolev space is defined by
[…]
    We also denote the norms
    $\|F\|_{k}\vcentcolon=\bigg{(}\int_{\Omega}|D^{k}F|^{2}{\mathrm{d}}x\bigg{)}^{\frac{1}{2}},\qquad\|F\|\vcentcolon=\|F\|_{0},\qquad\|F\|_{\infty}\vcentcolon=\sup\limits_{x\in\Omega}|F|.$
3 A-Priori Estimates
Define the difference functions:
$\phi\vcentcolon=\rho-\tilde{\rho},\qquad\psi\vcentcolon=u-\tilde{u},\qquad(\phi_{0},\psi_{0})\vcentcolon=(\phi,\psi)|_{t=0},$
and denote $D_{t}\vcentcolon=\partial_{t}+u\cdot\nabla$ as the material derivative. Then (1.1) is rewritten as
$\mathcal{L}^{0}(\phi,\psi)\vcentcolon=$ $D_{t}\phi+\rho_{+}\textnormal{{div}}\psi=f^{0},$ (3.8a)
$\mathcal{L}^{i}(\phi,\psi)\vcentcolon=$ $\partial_{t}\psi^{i}-\frac{\mu}{\rho_{+}}\Delta\psi^{i}-\frac{\mu+\lambda}{\rho_{+}}\partial_{x_{i}}\textnormal{{div}}\psi+q_{+}\partial_{x_{i}}\phi=f^{i},$ (3.8b)
for $i=1,2,3$, where $q(\rho)\vcentcolon=P^{\prime}(\rho)/\rho$ and $q_{+}\vcentcolon=q(\rho_{+})$, and
$f^{0}\vcentcolon=$ $-\phi\textnormal{{div}}\psi+(\rho_{+}-\tilde{\rho})\textnormal{{div}}\psi-\psi\cdot\nabla\tilde{\rho}-\phi\textnormal{{div}}\tilde{u},$ (3.8c)
$f^{i}\vcentcolon=$ $-\frac{\phi+\tilde{\rho}-\rho_{+}}{\rho_{+}(\phi+\tilde{\rho})}\big{\{}\mu\Delta\psi^{i}+(\mu+\lambda)\partial_{x_{i}}\textnormal{{div}}\psi\big{\}}+\tilde{f}^{i},$ (3.8d)
$\tilde{f}^{i}\vcentcolon=$ $-(\psi\cdot\nabla)\psi^{i}-(\tilde{u}\cdot\nabla)\psi^{i}-(\psi\cdot\nabla)\tilde{u}^{i}-\frac{\phi}{\phi+\tilde{\rho}}(\tilde{u}\cdot\nabla)\tilde{u}^{i}$
+$+\{q_{+}-q(\rho)\}\partial_{x_{i}}\phi-\frac{P^{\prime}(\rho_{+})-P^{\prime}(\rho)}{\phi+\tilde{\rho}}\partial_{x_{i}}\tilde{\rho},$ (3.8e)
for $i=1,2,3$. In addition, we denote that $\mathcal{L}\equiv\hat{e}_{1}\mathcal{L}^{1}+\hat{e}_{2}\mathcal{L}^{2}+\hat{e}_{3}\mathcal{L}^{3}$ as the vector-valued operator, and $f\equiv\hat{e}_{1}f^{1}+\hat{e}_{2}f^{2}+\hat{e}_{3}f^{3}$.
Moreover, we define the energy norm:
$N^{2}(T)\vcentcolon=$ $\sup\limits_{0\leq t\leq T}\big{\{}\|(\phi,\psi)(t,\cdot)\|_{3}^{2}+\|\partial_{t}\phi(t,\cdot)\|_{2}^{2}+\|\partial_{t}\psi(t,\cdot)\|_{1}^{2}\big{\}}$
$+\int_{0}^{T}\!\big{\{}\|\nabla\phi\|_{2}^{2}+\|\partial_{t}\phi\|_{3}^{2}+\|\nabla\psi\|_{3}^{2}+\|\partial_{t}\psi\|_{2}^{2}\big{\}}\,{\mathrm{d}}t.$ (3.9)
To perform the continuity argument using a-priori estimates, we need the existence and uniqueness of a local-in-time solution. This is stated in the following lemma.
Lemma 3.1.
Suppose the initial data $(\rho_{0},u_{0})$ satisfies $\|(\phi_{0},\psi_{0})\|_{3}<\infty$. Then there exists $T_{\ast}>0$, depending only on $\|(\phi_{0},\psi_{0})\|_{3}$, such that the initial-boundary value problem (3.8) and (1.2)–(1.4) has a unique solution $(\rho,u)$ satisfying $N(T)<\infty$ for $T\in[0,T_{\ast})$.
The proof of Lemma 3.1 is obtained using the same procedure of [14]. It is a standard argument for the local-in-time well-poseness theorem, and we omit the detail in the present paper.
Given initial norm $N(0)>0$. By continuity of $T\mapsto N(T)$, there exists a time $T_{0}>0$ such that,
$N(T)\leq 2N(0)\qquad\text{for all }\ T\in[0,T_{0}].$ (3.10)
Then by the Sobolev embedding theorem $H^{2}(\Omega)\xhookrightarrow{}\mathcal{C}^{0}(\Omega)$, there exists a positive constant $\varepsilon=\varepsilon(\rho_{+},u_{b},\gamma,K,\mu,\lambda)$ such that if $|u_{b}|+N(0)\leq\varepsilon$ then
$\frac{\rho_{+}}{2}\leq\rho(t,x)\leq\frac{3}{2}\rho_{+}\qquad\text{for }\ (t,x)\in[0,T_{0}]\times\Omega.$ (3.11)
Moreover, it also follows by the Sobolev embedding theorem that there exists a generic constant $C>0$ such that for all $T>0$,
$\sup\limits_{0\leq t\leq T}\big{\|}\big{(}\phi,\nabla\phi,\psi,\nabla\psi,\partial_{t}\phi\big{)}(t,\cdot)\big{\|}_{\infty}\leq CN(T),$ (3.12a)
$\int_{0}^{T}\big{\|}\big{(}\nabla\phi,\partial_{t}\phi,\nabla\partial_{t}\phi,\nabla^{2}\psi,\partial_{t}\psi\big{)}(t,\cdot)\big{\|}_{\infty}\,{\mathrm{d}}t\leq CN(T).$ (3.12b)
The main aim of this section is derive the following theorem:
Theorem 3.2.
Let $(\tilde{\rho},\tilde{u})$ be the stationary solution to (1.5). Suppose $(\rho,u)$ is the solution to the problem (1.1)–(1.4) with maximum time of existence $T_{\ast}>0$. There exist constants $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$, $\varepsilon=\varepsilon(\rho_{+},\gamma,K,\mu,\lambda)>0$, and $T_{0}=T_{0}(\varepsilon)\in(0,T_{\ast}]$ which are independent of $u_{b}$ and $(\rho_{0},u_{0})$ such that, if $|u_{b}|+N(0)\leq\varepsilon$, then
$N(T_{0})\leq CN(0).$
3.1 $L^{2}$ Energy Estimate
For $\zeta$, $\xi\geq 0$, we define the potential energy density $H(\zeta,\xi)$ to be
$H(\zeta,\xi)\!\vcentcolon=\!\zeta\!\int_{\xi}^{\zeta}\!\dfrac{P(z)-P(\xi)}{z^{2}}\,{\mathrm{d}}z\!=\!\left\{\begin{aligned} &\dfrac{K}{\gamma-1}\big{\{}\zeta^{\gamma}-\xi^{\gamma}-\gamma\xi^{\gamma-1}(\zeta-\xi)\big{\}}&&\text{if }\ \gamma>1,\\ &K\xi\big{(}1-\dfrac{\zeta}{\xi}+\dfrac{\zeta}{\xi}\ln\dfrac{\zeta}{\xi}\big{)}&&\text{if }\ \gamma=1.\end{aligned}\right.$ (3.13)
It can be verified that $H(\zeta,\xi)$ is a positive convex function achieving global minimum at $\{\zeta=\xi\}$ with $H(\zeta,\xi)=0$ if and only if $\zeta=\xi$. Moreover $H(\zeta,\xi)$ satisfies
$\zeta\partial_{\zeta}H(\zeta,\xi)=H(\zeta,\xi)+P(\zeta)-P(\xi),\qquad\xi\partial_{\xi}H(\zeta,\xi)=-P^{\prime}(\xi)(\zeta-\xi),$ (3.14a)
$\zeta\partial_{\zeta}H(\zeta,\xi)+\xi\partial_{\xi}H(\zeta,\xi)=\gamma H(\zeta,\xi),$ (3.14b)
$P(\zeta)-P(\xi)-P^{\prime}(\xi)(\zeta-\xi)=(\gamma-1)H(\zeta,\xi).$ (3.14c)
With $H(\cdot,\cdot)$, we define the $L^{2}$ relative energy density $\mathcal{E}$ to be:
:$\mathcal{E}\vcentcolon=\dfrac{1}{2}\rho|u-\tilde{u}|^{2}+H(\rho,\tilde{\rho}),\qquad\mathcal{E}_{0}\equiv\mathcal{E}|_{t=0}.$ (3.15)
Using Taylor’s Theorem, we obtain the following proposition,
Proposition 3.3.
If there exist constants $0<\rho_{\ast}<\rho^{\ast}<\infty$ such that $\rho_{\ast}\leq\rho,\,\tilde{\rho}\leq\rho^{\ast}$, then
$C^{-1}|\rho-\tilde{\rho}|^{2}\leq H(\rho,\tilde{\rho})\leq C|\rho-\tilde{\rho}|^{2},$
for some constant $C=C(\rho^{\ast},\rho_{\ast},\gamma,K)>0$.
Lemma 3.4.
There exists $\delta_{0}=\delta_{0}(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that if $|u_{b}|\leq\delta_{0}$ then,
$\sup\limits_{0\leq t\leq T}\int_{\Omega}\mathcal{E}(t,x)\,{\mathrm{d}}x+\iint_{Q_{T}}\big{\{}\dfrac{\mu}{2}|\nabla\psi|^{2}+(\mu+\lambda)|\textnormal{{{div}}}\psi|^{2}\big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$+|u_{b}|\int_{0}^{T}\!\!\!\!\int_{\{|y|=1\}}\!\!\!\!\!H(\rho,\tilde{\rho})\,{\mathrm{d}}\mathcal{S}_{y}{\mathrm{d}}t+|u_{b}|^{3}\iint_{Q_{T}}\dfrac{|\phi|^{2}}{|x|^{7}}\,{\mathrm{d}}x{\mathrm{d}}t+|u_{b}|\iint_{Q_{T}}\!\!\frac{|x\cdot\psi|^{2}}{|x|^{9}}\,{\mathrm{d}}x{\mathrm{d}}t\leq C\mathcal{E}_{0},$
for all $T\in[0,T_{0}]$, where $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ is some constant independent of $u_{b}$.
Proof.
Equations (1.1a) and (1.5a) yields that
$\partial_{t}H(\rho,\tilde{\rho})+\big{(}P(\rho)-P(\tilde{\rho})\big{)}\textnormal{{div}}\psi$
$=$ $(1-\gamma)H(\rho,\tilde{\rho})\textnormal{{div}}\tilde{u}-\textnormal{{div}}\big{(}uH(\rho,\tilde{\rho})\big{)}-\dfrac{\phi}{\tilde{\rho}}(\psi\cdot\nabla)P(\tilde{\rho}).$
Integrating the above in $x\in\Omega$, and using the boundary condition (1.2a), we obtain
$\dfrac{{\mathrm{d}}}{{\mathrm{d}}t}\int_{\Omega}\!H(\rho,\tilde{\rho})(t,x)\,{\mathrm{d}}x+\int_{\Omega}\!\!\big{(}P(\rho)-P(\tilde{\rho})\big{)}\textnormal{{div}}\Psi{\mathrm{d}}x-u_{b}\int_{\{|y|=1\}}\!\!\!\!H(\rho,\tilde{\rho})(t,y)\,{\mathrm{d}}S_{y}$
$=(1-\gamma)\int_{\Omega}\!\!H(\rho,\tilde{\rho})\textnormal{{div}}\tilde{u}\,{\mathrm{d}}x-\int_{\Omega}\dfrac{\phi}{\tilde{\rho}}(\psi\cdot\nabla)P(\tilde{\rho})\,{\mathrm{d}}x.$ (3.16)
Using (1.1) and (1.5), we have
$\partial_{t}\big{(}\rho\dfrac{|\psi|^{2}}{2}\big{)}-\big{(}P(\rho)-P(\tilde{\rho})\big{)}\textnormal{{div}}\psi+\mu|\nabla\psi|^{2}+(\mu+\lambda)|\textnormal{{div}}\psi|^{2}+\textnormal{{div}}\mathcal{F}$
$=$ $-\psi\cdot\big{\{}\phi(\psi\cdot\nabla)\tilde{u}+\phi(\tilde{u}\cdot\nabla)\tilde{u}+\tilde{\rho}(\psi\cdot\nabla)\tilde{u}\big{\}},$
where
:$\mathcal{F}\vcentcolon=\rho u\dfrac{|\psi|^{2}}{2}+\psi\big{(}P(\rho)-P(\tilde{\rho})\big{)}-\mu(\nabla\psi)^{\top}\cdot\psi-(\mu+\lambda)\psi\textnormal{{div}}\psi.$
Integrating the above in $x\in\Omega$, and using the boundary condition (1.2), we have
$\dfrac{{\mathrm{d}}}{{\mathrm{d}}t}\int_{\Omega}\!\!\rho\dfrac{|\psi|^{2}}{2}\,{\mathrm{d}}x-\int_{\Omega}\!\!\big{(}P(\rho)-P(\tilde{\rho})\big{)}\textnormal{{div}}\psi\,{\mathrm{d}}x+\mu\int_{\Omega}\!\!|\nabla\psi|^{2}\,{\mathrm{d}}x+(\mu+\lambda)\int_{\Omega}\!\!|\textnormal{{div}}\psi|^{2}\,{\mathrm{d}}x$
$=-\int_{\Omega}\!\!\psi\cdot\big{\{}\phi(\psi\cdot\nabla)\tilde{u}+\phi(\tilde{u}\cdot\nabla)\tilde{u}+\tilde{\rho}(\psi\cdot\nabla)\tilde{u}\big{\}}\,{\mathrm{d}}x.$ (3.17)
Next, by the steady state momentum equation 1.5b, it follows that
$\psi\cdot\big{\{}\phi(\psi\cdot\nabla)\tilde{u}+\phi(\tilde{u}\cdot\nabla)\tilde{u}+\tilde{\rho}(\psi\cdot\nabla)\tilde{u}\big{\}}+\dfrac{\phi}{\tilde{\rho}}(\psi\cdot\nabla)P(\tilde{\rho})=\rho\psi\cdot(\psi\cdot\nabla)\tilde{u}+\phi\psi\cdot\dfrac{L\tilde{u}}{\tilde{\rho}}.$
Combining (3.1) and (3.1), then using definition 3.15, we obtain
$\dfrac{{\mathrm{d}}}{{\mathrm{d}}t}\int_{\Omega}\!\mathcal{E}\,{\mathrm{d}}x+|u_{b}|\!\!\int_{\{|y|=1\}}\!\!\!H(\rho,\tilde{\rho})(t,y)\,{\mathrm{d}}\mathcal{S}_{y}+\int_{\Omega}\big{\{}\mu|\nabla\psi|^{2}+(\mu+\lambda)|\textnormal{{div}}\psi|^{2}\big{\}}\,{\mathrm{d}}x$
$=$ $(1-\gamma)\int_{\Omega}\!\!H(\rho,\tilde{\rho})\textnormal{{div}}\tilde{u}\,{\mathrm{d}}x-\int_{\Omega}\!\!\big{\{}\rho\psi^{\top}\cdot\nabla\tilde{u}\cdot\psi+\phi\psi\cdot\dfrac{L\tilde{u}}{\tilde{\rho}}\big{\}}\,{\mathrm{d}}x$ (3.18)
Since $\tilde{u}=\frac{x}{r}\tilde{U}(r)$, for $r=|x|$, it follows that $\nabla\tilde{u}$ is a symmetric matrix and
$\nabla\tilde{u}=(\nabla\tilde{u})_{+}+(\nabla\tilde{u})_{-}\quad\text{and}\quad\textnormal{{div}}\tilde{u}=\partial_{r}\tilde{U}(r)+(n-1)\dfrac{\tilde{U}(r)}{r},$
where
$(\nabla\tilde{u})_{+}\vcentcolon=\big{(}\partial_{r}\tilde{U}(r)-\dfrac{\tilde{U}(r)}{r}\big{)}\dfrac{x\otimes x}{r^{2}},\qquad(\nabla\tilde{u})_{-}\vcentcolon=\dfrac{\tilde{U}(r)}{r}\mathbb{I}_{3}.$
By Lemma 1.1, one can show that
$\textnormal{{div}}\tilde{u}=\dfrac{\partial_{r}(r^{2}\tilde{U})}{r^{2}}=-\tilde{\rho}(1)u_{b}\dfrac{\partial_{r}\tilde{\rho}(r)}{r^{2}|\tilde{\rho}(r)|^{2}}=\tilde{\rho}(1)|u_{b}|\dfrac{\partial_{r}\tilde{\rho}(r)}{r^{2}|\tilde{\rho}(r)|^{2}}\geq\frac{|u_{b}|}{Cr^{7}}.$ (3.19)
Moreover, using the above and Lemma 1.1, it also follows that for all $V\in\mathbb{R}^{3}$,
$V^{\top}\cdot(\nabla\tilde{u})_{+}\cdot V=$$\big{\{}\partial_{r}\tilde{U}-\dfrac{\tilde{U}}{r}\big{\}}\big{|}\dfrac{x}{r}\cdot V\big{|}^{2}$
$=$ $\big{\{}\textnormal{{div}}\tilde{u}+\dfrac{3|u_{b}|\tilde{\rho}(1)}{r^{3}\tilde{\rho}(r)}\big{\}}\big{|}\dfrac{x}{r}\cdot V\big{|}^{2}\geq\frac{|u_{b}|}{C}\frac{|x\cdot V|^{2}}{r^{9}}.$
Using this in (3.1), we obtain
$\dfrac{{\mathrm{d}}}{{\mathrm{d}}t}\int_{\Omega}\mathcal{E}\,{\mathrm{d}}x+|u_{b}|\!\!\int_{\{|y|=1\}}\!\!\!H(\rho,\tilde{\rho})(t,y)\,{\mathrm{d}}\mathcal{S}_{y}+\int_{\Omega}\big{\{}\mu|\nabla\psi|^{2}+(\mu+\lambda)|\textnormal{{div}}\psi|^{2}\big{\}}\,{\mathrm{d}}x$
$+(\gamma-1)\int_{\Omega}\!\!H(\rho,\tilde{\rho})\textnormal{{div}}\tilde{u}\,{\mathrm{d}}x+\frac{|u_{b}|}{C}\int_{\Omega}\frac{|x\cdot\psi|^{2}}{|x|^{9}}\,{\mathrm{d}}x$
$=$ $-\int_{\Omega}\!\!\rho\psi^{\top}\cdot(\nabla\tilde{u})_{-}\cdot\psi\,{\mathrm{d}}x-\int_{\Omega}\!\!\phi\psi\cdot\dfrac{L\tilde{u}}{\tilde{\rho}}\,{\mathrm{d}}x$ (3.20)
By Lemma 1.1, (3.11), Propositions 3.3 and A.1, we have
$\Big{|}\int_{\Omega}\rho\psi^{\top}\cdot(\nabla\tilde{u})_{-}\cdot\psi\,{\mathrm{d}}x\Big{|}\leq$ $\frac{3\rho_{+}|u_{b}|}{2}\int_{\Omega}\dfrac{|\psi|^{2}}{|x|^{3}}\,{\mathrm{d}}x$
$\leq$ $9\rho_{+}|u_{b}|\int_{\Omega}|\nabla\psi|^{2}\,{\mathrm{d}}x\leq\dfrac{\mu}{4}\int_{\Omega}|\nabla\psi|^{2}\,{\mathrm{d}}x,$ (3.21)
where we choose $|u_{b}|\leq\mu/(36\rho_{+})$. In addition, by Lemma 1.1,
$\textnormal{{div}}\tilde{u}=\tilde{\rho}(1)|u_{b}|\dfrac{\partial_{r}\tilde{\rho}}{r^{2}\tilde{\rho}^{2}}\geq\frac{|u_{b}|^{3}}{Cr^{7}},\quad|L\tilde{u}|=\mu\tilde{\rho}(1)|u_{b}|\cdot\Big{|}\dfrac{\partial_{r}^{2}\tilde{\rho}}{r^{2}\tilde{\rho}^{2}}-2\dfrac{\partial_{r}\tilde{\rho}}{r^{3}\tilde{\rho}^{2}}-2\dfrac{|\partial_{r}\tilde{\rho}|^{2}}{r^{2}\tilde{\rho}^{3}}\Big{|}\leq\frac{C|u_{b}|^{3}}{r^{8}}.$
Using the above, and Proposition 3.3, we obtain that
$(\gamma-1)\int_{\Omega}H(\rho,\tilde{\rho})\textnormal{{div}}\tilde{u}\,{\mathrm{d}}x\geq\frac{|u_{b}|^{3}}{C_{1}}\int_{\Omega}\dfrac{|\phi|^{2}}{|x|^{7}}\,{\mathrm{d}}x,$ (3.22)
for some constant $C_{1}=C_{1}(\rho_{+},\mu,\gamma,K)>0$. Moreover, by Lemma 1.1, Cauchy-Schwarz’s inequality, and Proposition A.1, we get
$\Big{|}\int_{\Omega}\!\!\phi\psi\cdot\dfrac{L\tilde{u}}{\tilde{\rho}}\,{\mathrm{d}}x\Big{|}\leq C|u_{b}|^{3}\int_{\Omega}\dfrac{|\phi\psi|}{|x|^{8}}\,{\mathrm{d}}x\leq\dfrac{\mu}{4}\int_{\Omega}\!\!|\nabla\psi|^{2}\,{\mathrm{d}}x+\frac{|u_{b}|^{3}}{C_{1}}\int_{\Omega}\dfrac{|\phi|^{2}}{|x|^{7}}\,{\mathrm{d}}x,$ (3.23)
where we choose $|u_{b}|$ small enough such that $C|u_{b}|^{3}\leq 1/C_{1}$. Substituting (3.1)–(3.23) into (3.1), we obtain
$\dfrac{{\mathrm{d}}}{{\mathrm{d}}t}\int_{\Omega}\!\!\mathcal{E}\,{\mathrm{d}}x+\int_{\Omega}\!\!\big{\{}\dfrac{\mu}{2}|\nabla\psi|^{2}+(\mu+\lambda)|\textnormal{{div}}\psi|^{2}\big{\}}\,{\mathrm{d}}x$
$+|u_{b}|\int_{\{|y|=1\}}\!\!\!H(\rho,\tilde{\rho})(t,y)\,{\mathrm{d}}\mathcal{S}_{y}+\dfrac{|u_{b}|^{3}}{2C_{1}}\int_{\Omega}\dfrac{|\phi|^{2}}{|x|^{7}}\,{\mathrm{d}}x+\frac{|u_{b}|}{C}\int_{\Omega}\frac{|x\cdot\psi|^{2}}{|x|^{9}}\,{\mathrm{d}}x\leq 0.$
Integrating the above in time, we obtain the desired estimate. $\square$
3.2 Estiamtes for Temporal Derivatives
First, we remark that the temporal derivative preserves the boundary condition (1.2), in the sense that
$\partial_{t}\psi(t,x)=\partial_{t}^{2}\psi(t,x)=0\qquad\text{for }\ (t,x)\in\Gamma_{T}\equiv[0,T]\times\partial\Omega.$ (3.24)
Lemma 3.5.
There exists a positive constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for all $T\in[0,T_{0}]$ and $k=0,1$,
$\sup\limits_{0\leq t\leq T}\big{\{}\|\partial_{t}(\phi,\psi)\|^{2}+\|\nabla\partial_{t}^{k}\psi\|^{2}\big{\}}+\!\!\int_{0}^{T}\!\!\!\!\big{\{}\|\nabla\partial_{t}\psi\|^{2}\!+\!\|\partial_{t}^{k+1}(\phi,\psi)\|^{2}\big{\}}{\mathrm{d}}t+\!\!\iint_{\Gamma_{T}}\!\!|\partial_{t}\phi|^{2}{\mathrm{d}}x{\mathrm{d}}t$
$\leq$ $C\big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\big{\}}.$
Proof.
Taking $\partial_{t}$ on (3.8a), then multiplying the resultant equality with $\frac{q_{+}}{\rho_{+}}\partial_{t}\phi$, we get
$\partial_{t}\big{(}\frac{q_{+}}{2\rho_{+}}|\partial_{t}\phi|^{2}\big{)}+\textnormal{{div}}\big{(}\frac{|\partial_{t}\phi|^{2}}{2}u\big{)}+q_{+}\partial_{t}\phi\textnormal{{div}}\partial_{t}\psi$
$=$ $\frac{q_{+}}{\rho_{+}}\big{\{}\partial_{t}\phi\partial_{t}f^{0}+\frac{|\partial_{t}\phi|^{2}}{2}\textnormal{{div}}u-\partial_{t}\phi(\partial_{t}\psi\cdot\nabla)\phi\big{\}}.$ (3.25)
In addition, taking $\partial_{t}$ on (3.8b), multiplying the resultant equality with $\partial_{t}\psi$, and integrating by parts, we get
$\partial_{t}\big{(}\frac{1}{2}|\partial_{t}\psi|^{2}\big{)}+\frac{\mu}{\rho_{+}}|\nabla\partial_{t}\psi|^{2}+\frac{\mu+\lambda}{\rho_{+}}|\textnormal{{div}}\partial_{t}\psi|^{2}-q_{+}\partial_{t}\phi\textnormal{{div}}\partial_{t}\psi$
$=$ $\partial_{t}\psi\cdot\partial_{t}f+\textnormal{{div}}\Big{\{}\frac{\mu}{\rho_{+}}\sum_{i=1}^{3}\partial_{t}\psi^{i}\nabla\partial_{t}\psi^{i}+\frac{\mu+\lambda}{\rho_{+}}\partial_{t}\psi\textnormal{{div}}\partial_{t}\psi-q_{+}\partial_{t}\psi\partial_{t}\phi\Big{\}}.$ (3.26)
Adding (3.25)–(3.26), integrating the resultant equality in $(t,x)\in Q_{T}$, applying the boundary condition (3.24), Lemma 1.1, and inequality (3.12), we obtain
$\frac{1}{2}\int_{\Omega}\!\!\big{\{}\frac{q_{+}}{\rho_{+}}|\partial_{t}\phi|^{2}+|\partial_{t}\psi|^{2}\big{\}}{\mathrm{d}}x\bigg{|}_{t=0}^{t=T}+\frac{|u_{b}|}{2}\iint_{\Gamma_{T}}\!\!|\partial_{t}\phi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t$
$+\iint_{Q_{T}}\!\!\Big{\{}\frac{\mu}{\rho_{+}}|\nabla\partial_{t}\psi|^{2}+\frac{\mu+\lambda}{\rho_{+}}|\textnormal{{div}}\partial_{t}\psi|^{2}\Big{\}}{\mathrm{d}}x{\mathrm{d}}t$
$=$ $\iint_{Q_{T}}\!\!\Big{\{}\frac{q_{+}}{\rho_{+}}\partial_{t}\phi\partial_{t}f^{0}+\partial_{t}\psi\cdot\partial_{t}f\Big{\}}{\mathrm{d}}x{\mathrm{d}}t+\frac{q_{+}}{\rho_{+}}\!\iint_{Q_{T}}\!\!\partial_{t}\phi\big{\{}\frac{\partial_{t}\phi}{2}\textnormal{{div}}u-(\partial_{t}\psi\cdot\nabla)\phi\big{\}}{\mathrm{d}}x{\mathrm{d}}t.$
$\leq$ $C\Big{\{}|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}.$ (3.27)
Next, taking $\partial_{t}^{k}$ on (3.8a), then multiplying the resultant equality with $\partial_{t}^{k+1}\phi$, we get
$|\partial_{t}^{k+1}\phi|^{2}=\partial_{t}^{k+1}\phi\big{\{}\partial_{t}f^{0}-\rho_{+}\textnormal{{div}}\partial_{t}^{k}\psi-\partial_{t}^{k}\big{(}u\cdot\nabla\phi\big{)}\big{\}}.$ (3.28)
Set $k=0,1$. Applying $\partial_{t}^{k}$ on (3.8b), then taking inner-product on the resultant equality with $\partial_{t}^{k+1}\psi$, we get
$\partial_{t}\Big{\{}\frac{\mu}{2\rho_{+}}|\nabla\partial_{t}^{k}\psi|^{2}+\frac{\mu+\lambda}{2\rho_{+}}\big{(}\textnormal{{div}}\partial_{t}^{k}\psi-\frac{\rho_{+}q_{+}}{\mu+\lambda}\partial_{t}^{k}\phi\big{)}^{2}-\frac{\rho_{+}q_{+}^{2}}{2(\mu+\lambda)}|\partial_{t}^{k}\phi|^{2}\Big{\}}+|\partial_{t}^{k+1}\psi|^{2}$
$=$ $\partial_{t}^{k+1}\psi\cdot\partial_{t}^{k}f-q_{+}\partial_{t}^{k+1}\phi\textnormal{{div}}\partial_{t}^{k}\psi$
$+\textnormal{{div}}\Big{\{}\frac{\mu}{\rho_{+}}\big{(}\nabla\partial_{t}^{k}\psi\big{)}^{\top}\cdot\partial_{t}^{k+1}\psi+\frac{\mu+\lambda}{\rho_{+}}\partial_{t}^{k+1}\psi\textnormal{{div}}\partial_{t}^{k}\psi-q_{+}\partial_{t}^{k}\phi\partial_{t}^{k+1}\psi\Big{\}}.$ (3.29)
Adding (3.28)–(3.29), integrating in $(t,x)\in Q_{T}$, using boundary condition (3.24) and inequality (3.12), we get
$\int_{\Omega}\!\!\Big{\{}\frac{\mu}{2\rho_{+}}|\nabla\partial_{t}^{k}\psi|^{2}+\frac{\mu+\lambda}{2\rho_{+}}\big{(}\textnormal{{div}}\partial_{t}^{k}\psi-\frac{\rho_{+}q_{+}}{\mu+\lambda}\partial_{t}^{k}\phi\big{)}^{2}\Big{\}}{\mathrm{d}}x\bigg{|}_{t=0}^{t=T}+\iint_{Q_{T}}\!\!\!\big{|}\partial_{t}^{k+1}(\phi,\psi)\big{|}^{2}\,{\mathrm{d}}x{\mathrm{d}}t$
$=$ $\frac{\rho_{+}q_{+}^{2}}{2(\mu+\lambda)}\|\partial_{t}^{k}\phi(t,\cdot)\|^{2}\big{|}_{t=0}^{t=T}+\iint_{Q_{T}}\partial_{t}^{k+1}\psi\cdot\partial_{t}^{k}f$
$+\iint_{Q_{T}}\!\!\!\partial_{t}^{k+1}\phi\Big{\{}\partial_{t}f^{0}-(\rho_{+}+q_{+})\textnormal{{div}}\partial_{t}^{k}\psi-\partial_{t}^{k}\big{(}u\cdot\nabla\phi\big{)}\Big{\}}{\mathrm{d}}x{\mathrm{d}}t,$
which holds true for $k=0,1$. Applying Lemma 3.4, inequality (3.12), and the previous estimate (3.27) on the above, we get
$\int_{\Omega}\!\!\Big{\{}\frac{\mu}{2\rho_{+}}|\nabla\partial_{t}^{k}\psi|^{2}+\frac{\mu+\lambda}{2\rho_{+}}\big{(}\textnormal{{div}}\partial_{t}^{k}\psi-\frac{\rho_{+}q_{+}}{\mu+\lambda}\partial_{t}^{k}\phi\big{)}^{2}\Big{\}}{\mathrm{d}}x+\frac{1}{2}\iint_{Q_{T}}\!\!\!\big{|}\partial_{t}^{k+1}(\phi,\psi)\big{|}^{2}\,{\mathrm{d}}x{\mathrm{d}}t$
$\leq$ $C\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}.$
This completes the proof. $\square$
3.3 Estimates for Angular Derivatives
Consider the spherical coordinate system $(r,\theta,\varphi)$, where $r\geq 1$ is the radial variable, $\theta\in[0,\pi]$ is the polar angle, and $\varphi\in[0,2\pi)$ is the azimuthal angle. For the angular derivatives, terms such as $1/\sin\theta$ occur, which form singularity at $\theta=0$ or $\pi$. To circumvent this issue, we consider the following partition of unity: let
$\Omega_{V}\vcentcolon=\Big{\{}x=(x_{1},x_{2},x_{3})^{\top}\in\mathbb{R}^{3}\ \Big{|}\ |x|\geq 1,\quad\arccos\big{(}\frac{x_{3}}{|x|}\big{)}\in\big{[}\frac{\pi}{9},\frac{8\pi}{9}\big{]}\Big{\}},$
$\Omega_{H}\vcentcolon=\Big{\{}x=(x_{1},x_{2},x_{3})^{\top}\in\mathbb{R}^{3}\ \Big{|}\ |x|\geq 1,\quad\arccos\big{(}\frac{x_{2}}{|x|}\big{)}\in\big{[}\frac{\pi}{9},\frac{8\pi}{9}\big{]}\Big{\}}.$
Then $\Omega=\Omega_{V}\cup\Omega_{H}$. On each domain $\Omega_{V}$ or $\Omega_{H}$, we define the spherical coordinate variables $(r,\theta_{V},\varphi_{V})$ and $(r,\theta_{H},\varphi_{H})$ as follows:
$x_{1}=r\cos\varphi_{V}\sin\theta_{V},\quad$ $x_{2}=r\sin\varphi_{V}\sin\theta_{V},\quad$ $x_{3}=r\cos\theta_{V},\quad$ for $\text{for }x\in\Omega_{V},$
$x_{1}=r\cos\varphi_{H}\sin\theta_{H},\quad$ $x_{2}=r\cos\theta_{H},\quad$ $x_{3}=r\sin\varphi_{H}\sin\theta_{H},\quad$ for $\text{for }x\in\Omega_{H},$
For $x\in\Omega_{V}$, we also define the spherical unit vectors
$\hat{r}\vcentcolon=\frac{x}{|x|},\quad\hat{\theta}_{V}\vcentcolon=\frac{x_{1}x_{3}\hat{e}_{1}+x_{2}x_{3}\hat{e}_{2}-(x_{1}^{2}+x_{2}^{2})\hat{e}_{3}}{|x|\sqrt{x_{1}^{2}+x_{2}^{2}}},\quad\hat{\varphi}_{V}\vcentcolon=\frac{-x_{2}\hat{e}_{1}+x_{1}\hat{e}_{2}}{\sqrt{x_{1}^{2}+x_{2}^{2}}}.$
For the case $x\in\Omega_{H}$, unit vectors $(\hat{r},\hat{\theta}_{H},\hat{\varphi}_{H})$ is defined by exchanging the terms $x_{2}\leftrightarrow x_{3}$ and $\hat{e}_{2}\leftrightarrow\hat{e}_{3}$ on the above expression. Let $\chi_{V}$, $\chi_{H}\in\mathcal{C}^{\infty}(\mathbb{R}^{3})$ be non-negative functions such that for all $x\in\mathbb{R}^{3}$, and $k=1,2,3,4$,
$0\leq\chi_{V},\,\chi_{H}\leq 1,\quad\chi_{V}+\chi_{V}\geq 1,\quad\text{supp}\chi_{V}\subseteq\Omega_{V},\quad\text{supp}\chi_{H}\subseteq\Omega_{H},$ (3.30a)
$|\nabla\chi_{H}|^{2}\leq C\chi_{H},\qquad|\nabla\chi_{V}|^{2}\leq C\chi_{V},\qquad|D^{k}\chi_{H}|+|D^{k}\chi_{V}|\leq C,$ (3.30b)
$\hat{r}\cdot\nabla\chi_{H}=\hat{\varphi}_{H}\cdot\nabla\chi_{H}=0,\qquad\hat{r}\cdot\nabla\chi_{V}=\hat{\varphi}_{V}\cdot\nabla\chi_{V}=0,$ (3.30c)
where $C>0$ is some generic constant. The detailed construction for $\chi_{V}$ and $\chi_{H}$ described above, can be found in Appendix B. Next, we give the precise definition for radial and angular derivatives. Consider the case $x\in\Omega_{V}$. We use the following conventions to denote the differential operators for spherical coordinate directions:
$\partial\in\{\partial_{\theta_{V}},\partial_{\varphi_{V}},\partial_{\theta_{H}},\partial_{\varphi_{H}}\},\qquad D\in\{\partial_{r},\partial_{\theta_{V}},\partial_{\varphi_{V}},\partial_{\theta_{H}},\partial_{\varphi_{H}}\},$
where
$\partial_{r}\vcentcolon=\frac{x}{|x|}\cdot\nabla,\qquad\partial_{\theta_{V}}\vcentcolon=|x|\hat{\theta}_{V}\cdot\nabla,\qquad\partial_{\varphi_{V}}\vcentcolon=\sqrt{x_{1}^{2}+x_{2}^{2}}\,\,\hat{\varphi}_{V}\cdot\nabla,$
$\partial_{\theta_{H}}\vcentcolon=|x|\hat{\theta}_{H}\cdot\nabla,\qquad\partial_{\varphi_{H}}\vcentcolon=\sqrt{x_{1}^{2}+x_{3}^{2}}\,\,\hat{\varphi}_{H}\cdot\nabla$
We remark that $\partial$ preserves boundary condition (1.2) in the sense that
$\chi\partial^{k}\psi|_{x\in\partial\Omega}=0\quad\text{for }\ \partial\in\{\partial_{\theta_{V}},\partial_{\varphi_{V}},\partial_{\theta_{H}},\partial_{\varphi_{H}}\}\ \text{ and }\ k=1,2,3.$ (3.31)
In what follows, we only derive estimates for one of $\Omega_{V}$ and $\Omega_{H}$, since the other case follows from the same argument. To simplify, we use the following notations,
$\chi\equiv\chi_{V}\ \text{or}\ \chi_{H},\qquad(\theta,\varphi)\equiv(\theta_{V},\varphi_{V})\ \text{or}\ (\theta_{H},\varphi_{H}),\qquad\tilde{\Omega}\equiv\Omega_{V}\ \text{or}\ \Omega_{H},$
$\tilde{Q}_{T}\equiv[0,T]\times\Omega_{V}\ \text{ or }\ [0,T]\times\Omega_{H},\qquad\tilde{\Gamma}_{T}\equiv[0,T]\times\partial\Omega_{V}\ \text{ or }\ [0,T]\times\partial\Omega_{H}.$
Lemma 3.6.
There exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for $\delta\in(0,1]$, $T\in[0,T_{0}]$, and $k=1,2,3$,
$\sup\limits_{0\leq t\leq T}\big{\|}\sqrt{\chi}(\partial^{k}\psi,\partial^{k}\phi)(t,\cdot)\big{\|}^{2}+\int_{0}^{T}\!\!\!\|\sqrt{\chi}D\partial^{k}\psi(t,\cdot)\|^{2}\,{\mathrm{d}}t$
$\leq$ $\frac{C}{\delta^{7}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+C\delta N^{2}(T).$
Proof.
Taking $\partial^{k}$ on equation (3.8a), and multiplying both sides of the resultant equality with $\chi\partial^{k}\phi$, we get
$\frac{q_{+}}{2\rho_{+}}\partial_{t}(\chi|\partial^{k}\phi|^{2})+q_{+}\chi\partial^{k}\phi\,\textnormal{{div}}\partial^{k}\psi=q_{+}\chi\partial^{k}\phi[\textnormal{{div}},\partial^{k}]\psi+\frac{q_{+}}{\rho_{+}}\chi\partial^{k}\phi\partial^{k}\big{\{}f^{0}-(u\cdot\nabla)\phi\big{\}}.$
Taking $\partial^{k}$ on equation (3.8b), and multiplying both sides of the resultant equality with $\chi\partial^{k}\psi$, we get
$\frac{1}{2}\partial_{t}(\chi|\partial^{k}\psi|^{2})-\frac{\mu}{\rho_{+}}\chi\partial^{k}\psi\cdot\Delta\partial^{k}\psi-\frac{\mu+\lambda}{\rho_{+}}\chi(\partial^{k}\psi\cdot\nabla)\textnormal{{div}}\partial^{k}\psi+q_{+}\chi(\partial^{k}\psi\cdot\nabla)\partial^{k}\phi$
$=$ $q_{+}\chi\partial^{k}\psi\cdot[\nabla,\partial^{k}]\phi+\frac{\mu}{\rho_{+}}\chi\partial^{k}\psi\cdot[\partial^{k},\Delta]\psi+\frac{\mu+\lambda}{\rho_{+}}\chi\partial^{k}\psi\cdot\big{[}\partial^{k},\nabla\textnormal{{div}}\big{]}\psi+\chi\partial^{k}\psi\cdot\partial^{k}f.$
Combining the two equations, integrating in $(t,x)\in Q_{T}\equiv[0,T]\times\Omega$, and integrating by parts using boundary condtion (3.31), we have
$\int_{\Omega}\frac{\chi}{2}\big{\{}|\partial^{k}\psi|^{2}+\frac{q_{+}}{\rho_{+}}|\partial^{k}\phi|^{2}\big{\}}(t,x)\,{\mathrm{d}}x\Big{|}_{t=0}^{t=T}$
$+\iint_{Q_{T}}\!\frac{\chi}{\rho_{+}}\Big{\{}\mu|\nabla\partial^{k}\psi|^{2}+(\mu+\lambda)|\textnormal{{div}}\partial^{k}\psi|^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$=$ $\iint_{Q_{T}}\!\!\Big{\{}q_{+}\partial^{k}\phi(\partial^{k}\psi\cdot\nabla)\chi-\frac{\mu}{\rho_{+}}\partial^{k}\psi^{\top}\cdot\nabla\partial^{k}\psi\cdot\nabla\chi-\frac{\mu+\lambda}{\rho_{+}}\textnormal{{div}}\partial^{k}\psi(\partial^{k}\psi\cdot\nabla)\chi\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$+\iint_{Q_{T}}\chi\Big{\{}q_{+}\partial^{k}\phi\cdot[\textnormal{{div}},\partial^{k}]\psi+q_{+}\partial^{k}\psi\cdot\big{[}\nabla,\partial^{k}\big{]}\phi\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$+\iint_{Q_{T}}\chi\Big{\{}\partial^{k}\psi\cdot\partial^{k}f+\frac{q_{+}}{\rho_{+}}\partial^{k}\phi\partial^{k}(f^{0}-u\cdot\nabla\phi)\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$+\frac{1}{\rho_{+}}\iint_{Q_{T}}\chi\partial^{k}\psi\cdot\Big{\{}\mu[\partial^{k},\Delta]\psi+(\mu+\lambda)\big{[}\partial^{k},\nabla\textnormal{{div}}\big{]}\psi\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t=\vcentcolon\sum_{i=1}^{4}\mathcal{I}_{i}.$ (3.32)
By the property of cut-off function $\chi$ in (3.30), we have for all $\epsilon\in(0,1]$,
$|\mathcal{I}_{1}|\leq$ $\iint_{Q_{T}}\!\!\frac{\chi}{4\rho_{+}}\Big{\{}\mu|\nabla\partial^{k}\psi|^{2}+(\mu+\lambda)|\textnormal{{div}}\partial^{k}\psi|^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t+\frac{\epsilon}{3}N^{2}(T)+\frac{C}{\epsilon}\iint_{\tilde{Q}_{T}}\!\!|\partial^{k}\psi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t.$
Using the commutation relations, Proposition B.1, we also get for all $\epsilon\in(0,1]$,
$|\mathcal{I}_{2}|\leq$ $\frac{\epsilon}{2}N^{2}(T)+\frac{C}{\epsilon}\sum_{m=0}^{k-1}\iint_{Q_{T}}\chi|\nabla\partial^{m}\psi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t.$
For the term $\mathcal{I}_{3}$, we recall the definition (3.8d). Then integrating by parts, applying Lemma 1.1 and inequality (3.12), we get for $k=1,2,3$,
$\bigg{|}\iint_{Q_{T}}\chi\partial^{k}\psi\cdot\partial^{k}(f-\tilde{f})\,\bigg{|}=\bigg{|}\sum_{i=1}^{3}\iint_{Q_{T}}\chi\partial^{k}\psi^{i}\partial^{k}\Big{(}\frac{\phi+\tilde{\rho}-\rho_{+}}{\rho_{+}(\phi+\tilde{\rho})}(L\psi)^{i}\Big{)}\bigg{|}$
$=$ $\bigg{|}\sum_{i=1}^{3}\!\iint_{Q_{T}}\!\!\!\big{\{}\partial\chi\partial^{k}\psi^{i}+\chi\partial^{k+1}\psi^{i}\big{\}}\partial^{k-1}\!\Big{(}\frac{\phi+\tilde{\rho}-\rho_{+}}{\rho_{+}(\phi+\tilde{\rho})}(L\psi)^{i}\Big{)}\bigg{|}\!\leq\!C\big{\{}|u_{b}|N^{2}(T)+N^{3}(T)\big{\}}.$
Thus by the above estimate, Lemma 1.1, and inequality (3.12), it follows that
$|\mathcal{I}_{3}|=$ $\bigg{|}\iint_{Q_{T}}\chi\partial^{k}\psi\cdot\big{\{}\partial^{k}f+\partial^{k}(f-\tilde{f})\big{\}}+\chi\frac{q_{+}}{\rho_{+}}\partial^{k}\phi\partial^{k}(f^{0}-u\cdot\nabla\phi){\mathrm{d}}x{\mathrm{d}}t\bigg{|}$
$\leq$ $CN^{3}(T)+C|u_{b}|N^{2}(T).$
Applying Propositions A.1 and B.1 we obtain that
$|\mathcal{I}_{4}|\leq$ $C\iint_{Q_{T}}\chi|\partial^{k}\psi|\Big{\{}\sum_{m=1}^{k+1}|\partial^{m}\psi|+\frac{|\psi|^{2}}{|x|^{4}}+\sum_{m=0}^{k-1}|D^{2}\partial^{m}\psi|\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$\leq$ $\frac{\epsilon}{3}N^{2}(T)+\frac{C}{\epsilon}\iint_{Q_{T}}\chi|\partial^{k}\psi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t.$
Substituting $\mathcal{I}_{1}$$\mathcal{I}_{4}$ into (3.32), it holds that for $k=1,2,3$, and $\epsilon\in(0,1]$,
$\int_{\Omega}\!\chi\big{\{}|\partial^{k}\psi|^{2}+\frac{q_{+}}{\rho_{+}}|\partial^{k}\phi|^{2}\big{\}}(T,x)\,{\mathrm{d}}x+\iint_{Q_{T}}\!\chi\Big{\{}\mu|\nabla\partial^{k}\psi|^{2}+(\mu+\lambda)|\textnormal{{div}}\partial^{k}\psi|^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$\leq$ $C\big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\big{\}}+\epsilon N^{2}(T)+\frac{C}{\epsilon}\sum_{m=0}^{k-1}\iint_{\tilde{Q}_{T}}|\nabla\partial^{m}\psi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t.$ (3.33)
Take $k=1$ and $\epsilon=\delta^{4}$ for some $\delta\in(0,1)$ in the above inequality. Then by Lemma 3.4,
$\int_{\Omega}\!\chi\big{\{}|\partial\psi|^{2}+|\partial\phi|^{2}\big{\}}(T,x)\,{\mathrm{d}}x+\iint_{Q_{T}}\!\chi\Big{\{}\mu|\nabla\partial\psi|^{2}+(\mu+\lambda)|\textnormal{{div}}\partial\psi|^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$\leq$ $\frac{C}{\delta^{4}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+\delta^{4}N^{2}(T).$ (3.34)
Next, take $k=2$ and $\epsilon=\delta^{2}$ in (3.33). Then applying Lemma 3.4, and substituting (3.34) into the resultant inequality, we get
$\int_{\Omega}\!\chi\big{\{}|\partial^{2}\psi|^{2}+|\partial^{2}\phi|^{2}\big{\}}(T,x)\,{\mathrm{d}}x+\iint_{Q_{T}}\!\!\chi\Big{\{}\mu|\nabla\partial^{2}\psi|^{2}+(\mu+\lambda)|\textnormal{{div}}\partial^{2}\psi|^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$\leq$ $\frac{C}{\delta^{6}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+C\delta^{2}N^{2}(T).$ (3.35)
Finally, take $k=3$ and $\epsilon=\delta$ in (3.33). Then applying Lemma 3.4, and substituting (3.34)–(3.35) into the resultant inequality, we get
$\int_{\Omega}\!\chi\big{\{}|\partial^{3}\psi|^{2}+|\partial^{3}\phi|^{2}\big{\}}(T,x)\,{\mathrm{d}}x+\iint_{Q_{T}}\!\!\chi\Big{\{}\mu|\nabla\partial^{3}\psi|^{2}+(\mu+\lambda)|\textnormal{{div}}\partial^{3}\psi|^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$\leq$ $\frac{C}{\delta^{7}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+C\delta N^{2}(T).$
This concludes the proof. $\square$
Corollary 3.7.
There exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for all $\delta\in(0,1]$, $T\in[0,T_{0}]$, and $k=0,1,2,3$,
$\iint_{Q_{T}}|D_{t}\phi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t\leq C\Big{\{}N^{2}(0)+N^{3}(T)+|u_{b}|N^{2}(T)\Big{\}},$
$\iint_{Q_{T}}\chi|\partial^{k}D_{t}\phi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t\leq\frac{C}{\delta^{7}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+C\delta N^{2}(T).$
Proof.
The equation (1.1a) yields $D_{t}\phi=-\rho_{+}\textnormal{{div}}\psi+f^{0}$. Integrating this equation in $(t,x)\in Q_{T}$, applying Lemmas 3.4 and inequality (3.12), we obtain the first estimate. Next, taking $\partial^{k}$ on (3.8a) yields $\partial^{k}D_{t}\phi=-\rho_{+}\textnormal{{div}}\partial^{k}\psi-\rho_{+}[\partial^{k},\textnormal{{div}}]\psi+\partial^{k}f^{0}$. Integrating this equation in $(t,x)\in Q_{T}$, applying Lemmas 3.4, 3.6, Proposition B.1, and inequality (3.12), we obtain the second estimate. $\square$
3.4 Damping Estimates for $\partial_{r}\phi$
Taking $\partial_{r}$ on (3.8a), we get
$-\rho_{+}\partial_{r}\textnormal{{div}}\psi=\partial_{t}\partial_{r}\phi+u\cdot\nabla\partial_{r}\phi-\partial_{r}f^{0}+[\partial_{r},(\psi\cdot\nabla)]\phi+[\partial_{r},(\tilde{u}\cdot\nabla)]\phi.$
Substituting this into the radial component of equation (3.8b), $\hat{r}\cdot\mathcal{L}(\phi,\psi)=\hat{r}\cdot f$, yields
$\partial_{r}D_{t}\phi+\beta\partial_{r}\phi=g_{1}\qquad\text{ or }\qquad D_{t}\partial_{r}\phi+\beta\partial_{r}\phi=g_{2},$ (3.36)
where $D_{t}\vcentcolon=\partial_{t}+(u\cdot\nabla)$, $\beta\vcentcolon=\rho_{+}^{2}q_{+}/(2\mu+\lambda)>0$, and
$g_{1}\vcentcolon=$ $\partial_{r}f^{0}+\frac{\beta}{q_{+}}\hat{r}\cdot(f-\partial_{t}\psi)+\frac{\mu\beta}{\rho_{+}q_{+}}\big{\{}\hat{r}\cdot\Delta\psi-\partial_{r}\textnormal{{div}}\psi\big{\}},$
$g_{2}\vcentcolon=$ $g_{1}+[(\psi\cdot\nabla),\partial_{r}]\phi+[(\tilde{u}\cdot\nabla),\partial_{r}]\phi.$
Lemma 3.8.
There exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for $T\in[0,T_{0}]$, and $k+l=0,1,2$,
$\sup\limits_{0\leq t\leq T}\big{\|}\sqrt{\chi}\partial_{r}^{l+1}\partial^{k}\phi\big{\|}^{2}+|u_{b}|\iint_{\tilde{\Gamma}_{T}}|\partial_{r}^{l+1}\partial^{k}\phi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t$
$+\int_{0}^{T}\!\Big{\{}\big{\|}\sqrt{\chi}\partial_{r}^{l+1}\partial^{k}\phi\big{\|}^{2}+\big{\|}\sqrt{\chi}\partial_{r}^{l+1}\partial^{k}D_{t}\phi\big{\|}^{2}\Big{\}}\,{\mathrm{d}}t$
$\leq$ $C\big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\big{\}}+C\!\!\int_{0}^{T}\!\!\!\Big{\{}\sum_{j=0}^{k+1}\big{\|}\sqrt{\chi}D\partial^{j}\psi\big{\|}_{l}^{2}+\big{\|}\sqrt{\chi}\partial_{r}^{l}\partial^{k}\partial_{t}\psi\big{\|}^{2}\Big{\}}\,{\mathrm{d}}t.$
Proof.
Taking derivative $\partial_{r}^{l}\partial^{k}$ on the first equation in (3.36), multiplying both sides by $\chi\rho\partial_{r}^{l+1}\partial^{k}D_{t}\phi$, and using continuity equation (1.1a), we get
$\partial_{t}\big{(}\frac{\beta}{2}\chi\rho|\partial_{r}^{l+1}\partial^{k}\phi|^{2}\big{)}+\textnormal{{div}}\big{(}\frac{\beta}{2}\chi\rho u|\partial_{r}^{l+1}\partial^{k}\phi|^{2}\big{)}+|\partial_{r}^{l+1}\partial^{k}D_{t}\phi|^{2}$
$=$ $\beta\chi\rho\partial_{r}^{l+1}\partial^{k}\phi[D_{t},\partial_{r}^{l+1}\partial^{k}]\phi+\chi\rho\partial_{r}^{l}\partial^{k}g_{1}\partial_{r}^{l+1}\partial^{k}D_{t}\phi+\beta\frac{\rho}{2}|\partial_{r}^{l+1}\partial^{k}\phi|^{2}(u\cdot\nabla)\chi.$
In addition, taking derivative $\partial_{r}^{l}\partial^{k}$ on the second equation in (3.36), multiplying both sides by $\chi\rho\partial_{r}^{l+1}\partial^{k}\phi$, and using continuity equation (1.1a), we get
$\partial_{t}\big{(}\frac{1}{2}\chi\rho|\partial_{r}^{l+1}\partial^{k}\phi|^{2}\big{)}+\textnormal{{div}}\big{(}\frac{1}{2}\chi\rho u|\partial_{r}^{l+1}\partial^{k}\phi|^{2}\big{)}+\beta\chi\rho\big{|}\partial_{r}^{l+1}\partial^{k}\phi\big{|}^{2}$
$=$ $\chi\rho\partial_{r}^{l+1}\partial^{k}\phi\big{\{}\partial_{r}^{l}\partial^{k}g_{2}+[D_{t}\partial_{r}\,,\partial_{r}^{l}\partial^{k}]\phi\big{\}}+\frac{\rho}{2}|\partial_{r}^{l+1}\partial^{k}\phi|^{2}(u\cdot\nabla)\chi.$
Adding the above two equalities together, integrating the resultant equality in $(t,x)\in[0,T]\times\Omega$, then it follows from boundary condition (1.2) and Cauchy-Schwarz’s inequality that
$\frac{\beta+1}{2}\int_{\Omega}\chi\rho|\partial_{r}^{l+1}\partial^{k}\phi|^{2}(t,x)\,{\mathrm{d}}x\Big{|}_{t=0}^{t=T}+\frac{\beta+1}{2}\int_{0}^{T}\!\!\!\!\int_{\partial\Omega}\!\!\!|u_{b}|\chi\rho|\partial_{r}^{l+1}\partial^{k}\phi|^{2}(t,y)\,{\mathrm{d}}y{\mathrm{d}}t$
$+\frac{1}{2}\iint_{Q_{T}}\chi\rho\Big{\{}|\partial_{r}^{l+1}\partial^{k}D_{t}\phi|^{2}+\beta|\partial_{r}^{l+1}\partial^{k}\phi|^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$\leq$ $C\iint_{Q_{T}}\chi\Big{\{}\big{|}[D_{t}\,,\partial_{r}^{l+1}\partial^{k}]\phi\big{|}^{2}+\big{|}[D_{t}\partial_{r}\,,\partial_{r}^{l}\partial^{k}]\phi\big{|}^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$+C\iint_{Q_{T}}\chi\Big{\{}\big{|}\partial_{r}^{l}\partial^{k}[(\psi\cdot\nabla),\partial_{r}]\phi\big{|}^{2}+\big{|}\partial_{r}^{l}\partial^{k}[(\tilde{u}\cdot\nabla),\partial_{r}]\phi\big{|}^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
$+C\!\!\iint_{Q_{T}}\!\!\chi\big{\{}|\partial_{r}^{l+1}\partial^{k}f^{0}|^{2}+|f|^{2}\big{\}}\,{\mathrm{d}}x{\mathrm{d}}t+C\!\!\iint_{Q_{T}}\!\!\chi\big{|}\partial_{r}^{l}\partial^{k}(\hat{r}\cdot\Delta\psi-\partial_{r}\textnormal{{div}}\psi)\big{|}^{2}\,{\mathrm{d}}x{\mathrm{d}}t$
$+C\iint_{Q_{T}}|\partial_{r}^{l+1}\partial^{k}\phi|^{2}|(u\cdot\nabla)\chi|\,{\mathrm{d}}x{\mathrm{d}}t+C\iint_{Q_{T}}\!\!\chi|\partial_{r}^{l}\partial^{k}\partial_{t}\psi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t=\vcentcolon\sum_{i=1}^{6}\mathcal{J}_{i}.$ (3.37)
Using the inequality (3.12), Proposition B.1, and the fact that $\partial_{t}$, $\partial_{r}$, $\partial$ commute with each other, we obtain
$|\mathcal{J}_{1}|=$ $\bigg{|}\iint_{Q_{T}}\chi\Big{\{}\big{(}[(u\cdot\nabla)\,,\partial_{r}^{l+1}\partial^{k}]\phi\big{)}^{2}+\big{(}[(u\cdot\nabla)\partial_{r}\,,\partial_{r}^{l}\partial^{k}]\phi\big{)}^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t\bigg{|}$
$\leq$ $C|u_{b}|^{2}N^{2}(T)+CN^{4}(T),$
$|\mathcal{J}_{2}|\leq$ $C|u_{b}|^{2}N^{2}(T)+N^{4}(T).$
By the definition (3.8c)–(3.8d), and inequality (3.12), we obtain
$|\mathcal{J}_{3}|=C\bigg{|}\iint_{Q_{T}}\chi\big{\{}|\partial_{r}^{l+1}\partial^{k}f^{0}|^{2}+|f|^{2}\big{\}}\,{\mathrm{d}}x{\mathrm{d}}t\bigg{|}\leq C|u_{b}|N^{2}(T)+CN^{3}(T).$
By Proposition B.2, it follows that
$|\mathcal{J}_{4}|=$ $C\bigg{|}\iint_{Q_{T}}\chi|\partial_{r}^{l}\partial^{k}(\hat{r}\cdot\Delta\psi-\partial_{r}\textnormal{{div}}\psi)|^{2}\,{\mathrm{d}}x{\mathrm{d}}t\bigg{|}\leq C\int_{0}^{T}\sum_{j=0}^{k+1}\big{\|}\sqrt{\chi}D\partial^{j}\psi\big{\|}_{l}^{2}\,{\mathrm{d}}t.$
Using Lemma 1.1, and inequality (3.12), we get
$|\mathcal{J}_{5}|=C\bigg{|}\iint_{Q_{T}}|\partial_{r}^{l+1}\partial^{k}\phi|^{2}|(\psi+\tilde{u})\cdot\nabla\chi|\,{\mathrm{d}}x\bigg{|}\leq C|u_{b}|N^{2}(T)+CN^{3}(T).$
Putting the estimates for $\mathcal{J}_{1}$$\mathcal{J}_{5}$ into (3.37), we obtain the desired estimate. $\square$
3.5 Estimates from elliptic and Stokes systems
Let $\nu_{1},\,\nu_{2}>0$. Given a vector-valued function :$\mathcal{R}\vcentcolon\Omega\to\mathbb{R}^{3}$, which belongs to some Sobolev space. Suppose $\zeta\vcentcolon\Omega\to\mathbb{R}^{3}$ is a solution to the elliptic system:
$-\nu_{1}\Delta\zeta^{i}-\nu_{2}\partial_{x_{i}}\textnormal{{div}}\zeta=$ $\mathcal{R}^{i}$ for $\text{for }\ x\in\Omega\ \text{ and }\ i=1,2,3,$ (3.38a)
$\zeta(x)=$ 00 0 for $\text{for }\ x\in\partial\Omega.$ (3.38b)
Lemma 3.9.
There exists a constant $C=C(\nu_{1},\nu_{2})>0$ such that for $k=2,3$,
$\|D^{k}\zeta\|\leq C\big{\{}\|\mathcal{R}\|_{k-2}+\|\zeta\|\big{\}}.$
The above estimate is well-known and its derivation is presented in [1]. The $L^{2}$ norm on $\zeta$ contained in the right hand side is due to the unboundedness of the exterior domain $\Omega$.
In addition, suppose $\alpha\vcentcolon\partial\Omega\to\mathbb{R}^{3}$, $h^{0}\vcentcolon\Omega\to\mathbb{R}$, and $h\vcentcolon\Omega\to\mathbb{R}^{3}$ are functions which belong to some Sobolev spaces. Let $w\vcentcolon\Omega\to\mathbb{R}$ and $v\vcentcolon\Omega\to\mathbb{R}^{3}$ be solution to the Stokes-type equations:
$\rho_{+}\textnormal{{div}}v$ $=h^{0}$ for $\text{for }\ x\in\Omega,$ (3.39a)
$-\mu\Delta v^{i}+\rho_{+}q_{+}\partial_{x_{i}}w$ $=h^{i}$ for $\text{for }\ x\in\Omega\ \text{ and }\ i=1,2,3,$ (3.39b)
$v(x)$ $=\alpha(x)$ for $\text{for }\ x\in\partial\Omega.$ (3.39c)
Lemma 3.10.
There exists $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for $k=2,3,4$,
$\|D^{k}v\|^{2}+\|D^{k-1}w\|^{2}\leq C\big{\{}\|h^{0}\|_{k-1}^{2}+\|h\|_{k-2}^{2}+\|\alpha\|_{H^{k-1/2}(\partial\Omega)}^{2}+\|Dv\|^{2}\big{\}}.$
The proof for the above lemma can be found in [15] (See Lemma 4.3). We also note that for the incompressible case: $h^{0}\equiv 0$, the system (3.39) reduces to the well-known Stokes equations, and the similar estimates can be found in [2, 5].
Lemma 3.11.
Let $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$ and $\partial\in\{\partial_{\theta},\partial_{\varphi}\}$. For $l=0,1,2$,
$\|D^{2+l}\psi\|^{2}+\|D^{l+1}\phi\|^{2}\leq C\Big{\{}\|\partial_{t}\psi\|_{l}^{2}+\|D_{t}\phi\|_{l+1}^{2}+\|D\psi\|^{2}+\|f^{0}\|_{l+1}^{2}+\|f\|_{l}^{2}\Big{\}}.$
For $k\geq 1$ and $k+l=1,2$, the following estimate holds
$\big{\|}\chi D^{2+l}\partial^{k}\psi\big{\|}^{2}+\big{\|}\chi D^{1+l}\partial^{k}\phi\big{\|}^{2}$
$\leq$ $C\Big{\{}\big{\|}\chi\partial^{k}D_{t}\phi\big{\|}_{l+1}^{2}+\big{\|}\chi\partial_{t}\partial^{k}\psi\big{\|}_{l}^{2}+\big{\|}f^{0}\big{\|}_{k+l+1}^{2}+\big{\|}f\big{\|}_{k+l}^{2}\Big{\}}$
$+C\Big{\{}\sum_{m=1}^{k}\big{\|}\partial^{m}\psi\big{\|}_{H^{l+1}(\tilde{\Omega})}^{2}+\sum_{m=0}^{k-1}\big{\|}D\partial^{m}\phi\big{\|}_{H^{l}(\tilde{\Omega})}^{2}\Big{\}}.$
Proof.
Applying the operator $\chi\partial^{k}$ on the equations (3.8a)–(3.8b), we have that $(v,w)\equiv(\chi\partial^{k}\phi,\chi\partial^{k}\psi)$ solves the Stokes-type system (3.39) with boundary condition $\chi\partial^{k}\psi|_{\partial\Omega}=0\equiv\alpha$, and
$h^{0}\vcentcolon=$ $\chi\partial^{k}f^{0}-\chi\partial^{k}D_{t}\phi+\rho_{+}\big{(}\partial^{k}\psi\cdot\nabla\big{)}\chi,$
$h^{i}\vcentcolon=$ $\rho_{+}\chi\partial^{k}f^{i}-\rho_{+}\chi\partial_{t}\partial^{k}\psi^{i}+\frac{\mu+\lambda}{\rho_{+}}\chi\partial^{k}\partial_{x_{i}}(f^{0}-D_{t}\phi)$
$+\mu\chi[\partial^{k},\Delta]\psi^{i}-\rho_{+}q_{+}\chi[\partial^{k},\partial_{x_{i}}]\phi+\mu\Delta\chi\partial^{k}\psi^{i}+2\mu(\nabla\chi\cdot\nabla)\partial^{k}\psi^{i},$
By (3.30), we have $|(\nabla\chi\cdot\nabla)F|\leq C\sqrt{\chi}|\partial_{\theta}F|$. Then by Lemma 3.11 and Propositions B.1, we have
$\big{\|}\chi D^{2+l}\partial^{k}\psi\big{\|}^{2}+\big{\|}\chi D^{1+l}\partial^{k}\phi\big{\|}^{2}$
$\leq$ $C\Big{\{}\big{\|}\chi\partial^{k}D_{t}\phi\big{\|}_{l+1}^{2}+\big{\|}\chi\partial_{t}\partial^{k}\psi\big{\|}_{l}^{2}+\big{\|}f^{0}\big{\|}_{k+l+1}^{2}+\big{\|}f\big{\|}_{k+l}^{2}\Big{\}}$
$+C\Big{\{}\sum_{m=1}^{k}\big{\|}\partial^{m}\psi\big{\|}_{H^{l+1}(\tilde{\Omega})}^{2}+\sum_{m=0}^{k-1}\big{\|}D\partial^{m}\phi\big{\|}_{H^{l}(\tilde{\Omega})}^{2}\Big{\}}$
This proves the second inequality. The first statement is shown by the exact same argument by setting $v=\psi$, $w=\phi$ in Lemma 3.10. $\square$
Lemma 3.12.
There exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for all $\delta\in(0,1]$ and $T\in[0,T_{0}]$,
$\sup\limits_{0\leq t\leq T}\|(\phi,\psi)\|_{2}^{2}+\int_{0}^{T}\!\!\big{\{}\|D\phi\|_{1}^{2}+\|(D\psi,\partial_{t}\phi)\|_{2}^{2}\big{\}}{\mathrm{d}}t$
$\leq$ $\frac{C}{\delta^{7}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+C\delta N^{2}(T).$
Proof.
Throughout this proof, we use the abbreviation:
$M_{\delta}(T)\vcentcolon=\frac{C}{\delta^{7}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+C\delta N^{2}(T).$ (3.40)
First, set $(k,l)=(0,0)$ and $(k,l)=(1,0)$ in Lemma 3.8. Then using Lemmas 3.4, 3.5, 3.6, and Corollary 3.7, we get
$\sup\limits_{0\leq t\leq T}\!\big{\|}\big{(}D\phi,\sqrt{\chi}D\partial\phi\big{)}\big{\|}^{2}\!+\!\!\int_{0}^{T}\!\!\!\big{\|}\big{(}\partial_{r}\phi,\sqrt{\chi}\partial_{r}\partial\phi,DD_{t}\phi,\sqrt{\chi}D\partial D_{t}\phi\big{)}\big{\|}^{2}{\mathrm{d}}t\leq M_{\delta}(T).$ (3.41)
Set $(k,l)=(0,0)$ in Lemma 3.11. Then by Lemma 3.5, Corollary 3.7, and (3.41),
$\int_{0}^{T}\!\!\big{\{}\big{\|}D^{2}\psi\big{\|}^{2}+\|D\phi\|^{2}\big{\}}{\mathrm{d}}t\leq M_{\delta}(T).$ (3.42)
Set $(k,l)=(1,0)$ in Lemma 3.11. Then by Lemmas 3.4, 3.5, 3.6, and (3.41)–(3.42),
$\int_{0}^{T}\!\!\big{\{}\big{\|}\chi D^{2}\partial\psi\big{\|}^{2}+\big{\|}\chi D\partial\phi\big{\|}^{2}\big{\}}{\mathrm{d}}t\leq M_{\delta}(T).$ (3.43)
Set $(k,l)=(0,1)$ in Lemma 3.8. Then using Lemmas 3.4, 3.5, 3.6, Corollary 3.7, and (3.42)–(3.43), we obtain
$\sup\limits_{0\leq t\leq T}\|D^{2}\phi\|^{2}+\int_{0}^{T}\!\!\Big{\{}\|\partial_{r}^{2}\phi\|^{2}+\|D^{2}D_{t}\phi\|^{2}\Big{\}}{\mathrm{d}}t\leq M_{\delta}(T).$ (3.44)
Set $(k,l)=(0,1)$ in Lemma 3.11. Then using Lemmas 3.4, 3.5, Corollary 3.7, inequalities (3.41) and (3.44), we have
$\int_{0}^{T}\big{\{}\big{\|}D^{3}\psi\big{\|}^{2}+\big{\|}D^{2}\phi\big{\|}^{2}\big{\}}{\mathrm{d}}t\leq M_{\delta}(T).$
Next, according to equation (3.8b), we set $\zeta\equiv\psi$ in Lemma 3.9. Then by Lemma 3.5, and (3.41), it follows that
$\sup\limits_{0\leq t\leq T}\|D^{2}\psi\|^{2}\leq C\sup\limits_{0\leq t\leq T}\Big{\{}\|\partial_{t}\psi\|^{2}+\|D\phi\|^{2}+\|f\|^{2}\Big{\}}\leq M_{\delta}(T).$
Finally, taking derivative $D$ on the equation (3.8a), we get $\partial_{t}D\phi=Df^{0}-D(u\cdot\nabla\phi)-\rho_{+}D\textnormal{{div}}\psi$. Using (3.12) and (3.42), we get that $\|\partial_{t}\phi\|_{L^{2}(0,T;H^{1}(\Omega))}\leq M_{\delta}(T)$. $\square$
Lemma 3.13.
There exists a constant $C\!=\!C(\rho_{+},\gamma,K,\mu,\lambda)>\!0$ such that for $\delta\in(0,1]$ and $T\in[0,T_{0}]$,
$\sup\limits_{0\leq t\leq T}\big{\|}(D^{3}\phi,D^{3}\psi)\big{\|}^{2}+\int_{0}^{T}\!\!\Big{\{}\big{\|}\big{(}D^{3}\phi,D^{3}\partial_{t}\phi\big{)}\big{\|}^{2}+\|D^{4}\psi\|^{2}\Big{\}}{\mathrm{d}}t$
$\leq$ $\frac{C}{\delta^{7}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+C\delta N^{2}(T).$
Proof.
For simplicity, we use the same notation $M_{\delta}(T)$ as defined in (3.40). Taking $\partial_{t}$ on the equation (3.8b), we verify that $\zeta\equiv\partial_{t}\psi$ solves the elliptic system (3.38), with $\nu_{1}\equiv\mu$, $\nu_{2}\equiv\mu+\lambda$ and
$\mathcal{R}^{i}\equiv\rho_{+}\partial_{t}f^{i}-\rho_{+}\partial_{t}^{2}\psi^{i}-\rho_{+}(\partial_{t}\psi\cdot\nabla)\psi^{i}-\rho_{+}(\psi\cdot\nabla)\partial_{t}\psi^{i}-\rho_{+}(\tilde{u}\cdot\nabla)\partial_{t}\psi^{i}-q_{+}\partial_{x_{i}}\partial_{t}\phi.$
Thus applying Lemma 3.9 on $\zeta\equiv\partial_{t}\psi$ with $k=0$, then using inequality (3.12), and Lemmas 3.5, 3.12, we obtain that
$\int_{0}^{T}\!\!\|D^{2}\partial_{t}\psi\|^{2}\,{\mathrm{d}}t\leq C\int_{0}^{T}\!\!\big{\{}\|\mathcal{R}\|^{2}+\|\partial_{t}\psi\|^{2}\big{\}}\,{\mathrm{d}}t\leq M_{\delta}(T).$ (3.45)
Take $(k,l)=(2,0)$ in Lemma 3.8. Then by Lemmas 3.4, 3.6, Corollary 3.7, and (3.45),
$\sup\limits_{0\leq t\leq T}\|\sqrt{\chi}D\partial^{2}\phi\|^{2}+\int_{0}^{T}\!\!\!\big{\{}\|\sqrt{\chi}\partial_{r}\partial^{2}\phi\|^{2}+\|\sqrt{\chi}D^{3}D_{t}\phi\|^{2}\big{\}}{\mathrm{d}}t\leq M_{\delta}(T).$ (3.46)
Take $(k,l)=(2,0)$ in Lemma 3.11. Then by Lemmas 3.6, 3.12 and (3.45)–(3.46),
$\int_{0}^{T}\!\!\big{\{}\|\chi D^{2}\partial^{2}\psi\|^{2}+\|\chi D\partial^{2}\phi\|^{2}\big{\}}{\mathrm{d}}t\leq M_{\delta}(T).$ (3.47)
Take $(k,l)=(1,1)$ in Lemma 3.8. Then by Lemmas 3.6, 3.12, Corollary 3.7, inequalities (3.45), and (3.47), we have
$\sup\limits_{0\leq t\leq T}\|\sqrt{\chi}D^{2}\partial\phi\|^{2}+\int_{0}^{T}\!\!\Big{\{}\big{\|}\sqrt{\chi}D^{2}\partial\phi\big{\|}^{2}+\big{\|}\sqrt{\chi}D^{2}\partial D_{t}\phi\big{\|}^{2}\Big{\}}{\mathrm{d}}t\leq M_{\delta}(T).$ (3.48)
Take $(k,l)=(1,1)$ in Lemma 3.11. Then by Lemma 3.12, inequalities (3.45) and (3.48),
$\int_{0}^{T}\!\!\Big{\{}\big{\|}\chi D^{3}\partial\psi\big{\|}^{2}+\big{\|}\chi D^{2}\partial\phi\big{\|}^{2}\Big{\}}{\mathrm{d}}t\leq M_{\delta}(T).$ (3.49)
Take $(k,l)=(0,2)$ in Lemma 3.8. Then by Lemmas 3.6, 3.12, Corollary 3.7, inequalities (3.45) and (3.49), we obtain
$\sup\limits_{0\leq t\leq T}\big{\|}D^{3}\phi\|^{2}+\int_{0}^{T}\!\!\big{\{}\|D^{3}\phi\|^{2}+\|D^{3}D_{t}\phi\|^{2}\big{\}}{\mathrm{d}}t\leq M_{\delta}(T).$ (3.50)
Take $(k,l)=(0,2)$ in Lemma 3.11. Then by Lemma 3.12, inequalities (3.45) and (3.50),
$\int_{0}^{T}\!\!\big{\{}\|D^{4}\psi\|^{2}+\|D^{3}\phi\|^{2}\big{\}}{\mathrm{d}}t\leq M_{\delta}(T).$ (3.51)
Finally, from the momentum equation (3.8b), we see that $\zeta\equiv\psi$ solves system (3.38) with $\nu_{1}\equiv\mu$, $\nu_{2}\equiv\mu+\lambda$, and $\mathcal{R}\equiv\rho_{+}\{f-\partial_{t}\psi-(u\cdot\nabla)\psi-q_{+}\nabla\phi\}$. Applying Lemma 3.9 on $\zeta\equiv\psi$ with $k=3$, we get $\sup_{0\leq t\leq T}\|D^{3}\psi(t,\cdot)\|^{2}\leq M_{\delta}(T)$. $\square$
Proof of Theorem 3.2.
By (3.9) and Lemmas 3.123.13, there exists a positive constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)$ such that for all $\delta\in(0,1]$ and $T\in[0,T_{0}]$,
$N^{2}(T)\leq\frac{C}{\delta^{7}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+C\delta N^{2}(T).$ (3.52)
Let $\delta=\frac{1}{4C}$, and set $\varepsilon=4^{-9}C^{-8}$. (3.10) and (3.52) implies that if $|u_{b}|+N(0)\leq\varepsilon$, then
$N^{2}(T)\leq 4^{7}C^{8}N^{2}(0)+\frac{3}{4}N^{2}(T)\qquad\text{ for all }\ T\in[0,T_{0}].$
Therefore $N^{2}(T_{0})\leq 4^{8}C^{8}N^{2}(0)$, and we conclude the proof of Theorem 3.2. $\square$
Appendix A Hardy’s Inequality
Proposition A.1.
Let $\Omega\vcentcolon=\{x\in\mathbb{R}^{n}\,|\,|x|\geq 1\}$ with $n\geq 3$. If $u\in H^{1}(\Omega)$, then $\frac{u}{|x|}\in L^{2}(\Omega)$, and the following estimate holds:
$\int_{\Omega}\dfrac{|u|^{2}}{|x|^{2}}\,{\mathrm{d}}x+\int_{\{|x|=1\}}\!\!|u|^{2}\,{\mathrm{d}}S_{x}\leq 2n\int_{\Omega}|\nabla u|^{2}\,{\mathrm{d}}x.$
Proof.
Fix $R>1$, and let $u\in(\mathcal{C}^{\infty}\cap H^{1})(\Omega)$. Since $\nabla(\frac{1}{|x|})=-\frac{x}{|x|^{3}}$, we have
$\int_{\Omega_{R}}\dfrac{|u|^{2}}{|x|^{2}}\,{\mathrm{d}}x=-\int_{\Omega_{R}}|u|^{2}\dfrac{x}{|x|}\cdot\nabla\big{(}\dfrac{1}{|x|}\big{)}\,{\mathrm{d}}x$
$=$ $\int_{\Omega_{R}}\!\!\Big{\{}2u\cdot\nabla u\cdot\dfrac{x}{|x|^{2}}+(n-1)\dfrac{|u|^{2}}{|x|^{2}}\Big{\}}\,{\mathrm{d}}x-\dfrac{1}{R}\int_{\{|x|=R\}}\!\!|u|^{2}\,{\mathrm{d}}S_{x}+\int_{\{|x|=1\}}\!\!|u|^{2}\,{\mathrm{d}}S_{x}.$
Thus reorganising the equation we obtain
$(n-2)\int_{\Omega_{R}}\dfrac{|u|^{2}}{|x|^{2}}\,{\mathrm{d}}x+\int_{\{|x|=1\}}\!\!|u|^{2}\,{\mathrm{d}}S_{x}=-2\int_{\Omega_{R}}u\cdot\nabla u\cdot\dfrac{x}{|x|^{2}}\,{\mathrm{d}}x+\dfrac{1}{R}\int_{\{|x|=R\}}\!\!|u|^{2}\,{\mathrm{d}}S_{x}.$
By Cauchy-Schwartz inequality, it follows that
$(n-2)\!\!\int_{\Omega_{R}}\dfrac{|u|^{2}}{|x|^{2}}\,{\mathrm{d}}x+\int_{\{|x|=1\}}\!\!|u|^{2}\,{\mathrm{d}}S_{x}$
$\leq$ $\dfrac{n-2}{2}\int_{\Omega_{R}}\!\!\dfrac{|u|^{2}}{|x|^{2}}\,{\mathrm{d}}x+\dfrac{2}{n-2}\int_{\Omega_{R}}\!\!|\nabla u|^{2}\,{\mathrm{d}}x+\int_{\{|x|=R\}}\!\!\dfrac{|u|^{2}}{R}\,{\mathrm{d}}S_{x}$ (A.53)
In addition, by divergence theorem and Cauchy-Schwartz inequality, we also have
$\int_{\{|x|=R\}}\!\!\dfrac{|u|^{2}}{R}\,{\mathrm{d}}S_{x}-\int_{\{|x|=1\}}\!\!\dfrac{|u|^{2}}{R^{2}}\,{\mathrm{d}}S_{x}=$ $\dfrac{1}{R^{2}}\int_{\Omega_{R}}\!\!\textnormal{{div}}(x|u|^{2})\,{\mathrm{d}}x=\dfrac{1}{R^{2}}\int_{\Omega_{R}}\!\!\big{\{}n|u|^{2}+2xu\nabla u\big{\}}\,{\mathrm{d}}x$
$\leq$ $\dfrac{n+1}{R^{2}}\int_{\Omega_{R}}\!\!|u|^{2}\,{\mathrm{d}}x+\int_{\Omega_{R}}\!\!|\nabla u|^{2}\,{\mathrm{d}}x.$
Substituting the above into (A), it follows that
$\int_{\Omega_{R}}\dfrac{|u|^{2}}{|x|^{2}}\,{\mathrm{d}}x+\dfrac{R^{2}-1}{R^{2}}\int_{\{|x|=1\}}\!\!|u|^{2}\,{\mathrm{d}}S_{x}\leq 2n\int_{\Omega_{R}}\!\!|\nabla u|^{2}\,{\mathrm{d}}x+\dfrac{2(n+1)}{n-2}\int_{\Omega_{R}}\!\!\dfrac{|u|^{2}}{R^{2}}\,{\mathrm{d}}x.$ (A.54)
By density argument, (A.54) holds for any $u\in H^{1}(\Omega)$. We remark that the trace of $u$ on the boundary domain $\{|x|=1\}$ is well defined by the trace embedding theorem, $\textrm{tr}\vcentcolon H^{1}(\Omega)\to L^{2}(\partial\Omega)$. Taking $R\to\infty$ on the resultant inequality, we obtain that
$\int_{\Omega}\dfrac{|u|^{2}}{|x|^{2}}\,{\mathrm{d}}x+\int_{\{|x|=1\}}\!\!|u|^{2}\,{\mathrm{d}}S_{x}\leq 2n\int_{\Omega}\!\!|\nabla u|^{2}\,{\mathrm{d}}x.$
This concludes the proof. $\square$
Appendix B Spherical Differential Operators
In this appendix, we study the properties of differential operators in spherical coordinate. First, we give the detailed construction of cut-off functions claimed in (3.30). Define the function $\tilde{\xi}(\theta)\in\mathcal{C}^{1}[0,\pi]$ by
$\tilde{\xi}(\theta)\vcentcolon=\left\{\begin{aligned} &0&&\text{if }\ \theta\in[0,\tfrac{\pi}{8})\\ &\tfrac{32}{\pi^{2}}(\theta-\tfrac{\pi}{8})^{2}&&\text{if }\ \theta\in[\tfrac{\pi}{8},\tfrac{\pi}{4}),\\ &1-\tfrac{32}{\pi^{2}}(\theta-\tfrac{3\pi}{8})^{2}&&\text{if }\ \theta\in[\tfrac{\pi}{4},\tfrac{3\pi}{8}),\\ &1&&\text{if }\ \theta\in[\tfrac{3\pi}{8},\tfrac{5\pi}{8}),\end{aligned}\right.\quad\left\{\begin{aligned} &1&&\text{if }\ \theta\in[\tfrac{3\pi}{8},\tfrac{5\pi}{8}),\\ &1-\tfrac{32}{\pi^{2}}(\theta-\tfrac{5\pi}{8})^{2}&&\text{if }\ \theta\in[\tfrac{5\pi}{8},\tfrac{3\pi}{4}),\\ &\tfrac{32}{\pi^{2}}(\theta-\tfrac{7\pi}{8})^{2}&&\text{if }\ \theta\in[\tfrac{3\pi}{4},\tfrac{7\pi}{8}),\\ &0&&\text{if }\ \theta\in[\tfrac{7\pi}{8},\pi].\end{aligned}\right.$
It can be verified that for all $\theta\in[0,\pi]$,
$0\leq\tilde{\xi}(\theta)\leq 1,\qquad|\tilde{\xi}^{\prime}(\theta)|^{2}\leq\frac{128}{\pi^{2}}\xi(\theta),\qquad\text{supp}\big{(}\tilde{\xi}\big{)}=\big{[}\frac{\pi}{8},\frac{7\pi}{8}\big{]}.$ (B.55)
Using the above properties, one can choose a suitable mollifier $\eta\in\mathcal{C}^{\infty}_{c}[-1,1]$ such that the convolution $\xi\vcentcolon=\tilde{\xi}\ast\eta$ satisfies,
$0\leq\xi\leq 1,\qquad|\xi^{\prime}|^{2}\leq C\xi,\qquad\text{supp}(\xi)\subseteq\big{[}\frac{\pi}{9},\frac{8\pi}{9}\big{]},\qquad\sup\limits_{\theta\in[0,\pi]}\big{|}\xi^{(k)}(\theta)\big{|}\leq C,$
for all $k=1,2,3,4$, where $C>0$ is some generic constant. Using this, we define the partition of unity functions as
$\chi_{V}(x)\vcentcolon=\xi\Big{(}\arccos\big{(}\frac{x_{3}}{|x|}\big{)}\Big{)}\quad\text{ and }\quad\chi_{H}(x)\vcentcolon=\xi\Big{(}\arccos\big{(}\frac{x_{2}}{|x|}\big{)}\Big{)}.$
Since $\text{supp}(\xi)=[\frac{\pi}{9},\frac{8\pi}{9}]$, one sees that $\chi_{V}=0$ for $\arccos(\frac{x_{3}}{|x|})\in[0,\frac{\pi}{9}]\cup[\frac{8\pi}{9},\pi]$. The gradient of $\chi_{V}$ is given by:
$\nabla\chi_{V}(x)=\xi^{\prime}\Big{(}\arccos\big{(}\frac{x_{3}}{|x|}\big{)}\Big{)}\frac{\hat{e}_{1}x_{1}x_{3}+\hat{e}_{2}x_{2}x_{3}-\hat{e}_{3}(x_{1}^{2}+x_{2}^{2})}{|x|\sqrt{x_{1}^{2}+x_{2}^{2}}}=\hat{\theta}_{V}\,\xi^{\prime}(\theta_{V}).$
Thus using (B.55), we also have for all $x\in\Omega$,
$|\nabla\chi_{V}(x)|^{2}\leq C\chi_{V}(x),\qquad\hat{r}\cdot\nabla\chi_{V}=\hat{\varphi}_{V}\cdot\nabla\chi_{V}=0.$
The statements for $\chi_{H}(x)$ is shown following the exact same argument. This proves the existence of functions described in (3.30). In what follows, we use for simplicity, the notations $(\theta,\varphi)\equiv(\theta_{V},\varphi_{V})$ or $(\theta_{H},\varphi_{H})$ and $\chi\equiv\chi_{V}$ or $\chi_{H}$.
Let $(r,\theta,\varphi)\in(0,\infty)\times[0,\pi]\times[0,2\pi)$ be the spherical coordinate. Then the following differential relations for unit basis vectors holds:
$\partial_{r}\hat{r}=0,\qquad$ $\partial_{\theta}\hat{r}=\hat{\theta},\qquad$ $\partial_{\varphi}\hat{r}=\hat{\varphi}\sin\theta,$ (B.56a)
$\partial_{r}\hat{\theta}=0,$ $\partial_{\theta}\hat{\theta}=-\hat{r},$ $\partial_{\varphi}\hat{\theta}=\hat{\varphi}\cos\theta,$ (B.56b)
$\partial_{r}\hat{\varphi}=0,$ $\partial_{\theta}\hat{\varphi}=0,$ $\partial_{\varphi}\hat{\varphi}=-\hat{r}\sin\theta-\hat{\theta}\cos\theta.$ (B.56c)
Let $F$ be a scalar-valued function and $V$ be a vector-valued function. Then the gradient, divergence, and Laplacian under spherical coordinate are given by
$\nabla F=\hat{r}\partial_{r}F+\hat{\theta}\frac{1}{r}\partial_{\theta}F+\hat{\varphi}\frac{1}{r\sin\theta}\partial_{\varphi}F,$ (B.57a)
$\textnormal{{div}}V=\frac{1}{r^{2}}\partial_{r}\big{(}r^{2}(\hat{r}\cdot V)\big{)}+\frac{1}{r\sin\theta}\partial_{\theta}\big{(}(\hat{\theta}\cdot V)\sin\theta\big{)}+\frac{1}{r\sin\theta}\partial_{\varphi}\big{(}\hat{\varphi}\cdot V\big{)},$ (B.57b)
$\Delta F=\frac{1}{r^{2}}\partial_{r}\big{(}r^{2}\partial_{r}F\big{)}+\frac{1}{r^{2}\sin\theta}\partial_{\theta}\big{(}\partial_{\theta}F\sin\theta\big{)}+\frac{1}{r^{2}\sin^{2}\theta}\partial_{\varphi}^{2}F.$ (B.57c)
Proposition B.1.
Let $\partial\in\{\partial_{\theta},\partial_{\varphi}\}$ and $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$. For given $N\in\mathbb{N}$, there exists a constant $C=C(N)>0$ such that for all scalar functions $F(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}$,
$\chi\big{|}[\partial^{N},\nabla]F\big{|}\leq C\chi\sum_{m=0}^{N-1}|\nabla\partial^{m}F|,\qquad\chi\big{|}[\partial^{N},\Delta]F\big{|}\leq C\chi\frac{1}{r^{2}}\sum_{m=1}^{N+1}|\partial^{m}F|.$
Moreover, for all $F(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}$ and $V(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$,
$\chi\big{|}[D^{N},(V\cdot\nabla)]F\big{|}\leq$ $C\chi\sum_{k=1}^{N}|D^{k}F|\sum_{m=0}^{N-k+1}|D^{m}V|,$
$\chi\big{|}[\partial^{N},\nabla\textnormal{{div}}]V\big{|}\leq$ $C\chi\frac{|V|}{|x|^{2}}+C\chi\sum_{m=0}^{N-1}|D^{2}\partial^{m}V|.$
Proof.
By product rule, binomial formula, and (B.57a), it holds that
$[\partial^{N},\nabla]F=\sum_{m=0}^{N-1}\binom{N}{m}\Big{\{}\big{(}\partial^{N-m}\hat{r}\big{)}\partial_{r}\partial^{m}F+\frac{\partial^{N-m}\hat{\theta}}{r}\partial_{\theta}\partial^{m}F+\partial^{N-m}\Big{(}\frac{\hat{\varphi}}{r\sin\theta}\Big{)}\partial_{\varphi}\partial^{m}F\Big{\}}.$
Thus using the differential relations (B.56), we obtain the first inequality. Next, by a similar calculation, we also obtain from (B.57c) that
$[\partial^{N},\Delta]F=\frac{1}{r^{2}}\sum_{m=0}^{N-1}\binom{N}{m}\Big{\{}\partial_{\theta}\partial^{m}F\partial^{N-m}\Big{(}\frac{\cos\theta}{\sin\theta}\Big{)}+\partial_{\varphi}^{2}\partial^{m}F\partial^{N-m}\Big{(}\frac{1}{\sin^{2}\theta}\Big{)}\Big{\}}.$
This yields the second inequality of the proposition. Let $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$, and set $V\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$ be a vector-valued function. We have by binomial formula and (B.57)
$\big{[}D^{N},(V\cdot\nabla)\big{]}F$
$=$ $\sum_{k=0}^{N-1}\binom{N}{k}D^{N-k}V\cdot\nabla D^{k}F+\sum_{m=0}^{N-1}\sum_{k=m+1}^{N}\binom{N}{k}\binom{k}{m}D^{N-k}V\cdot D^{k-m}\hat{r}D^{m}\partial_{r}F$
$+\sum_{m=0}^{N-1}\sum_{k=m+1}^{N}\binom{N}{k}\binom{k}{m}D^{N-k}V\cdot\Big{\{}D^{k-m}\Big{(}\frac{\hat{\theta}}{r}\Big{)}D^{m}\partial_{\theta}F+D^{k-m}\Big{(}\frac{\hat{\varphi}}{r\sin\theta}\Big{)}D^{m}\partial_{\varphi}F\Big{\}}.$
Multiplying both sides by $\chi$, then taking absolute value, we get
$\big{|}[D^{N},(V\cdot\nabla)]F\big{|}\leq$ $C\chi\Big{\{}\sum_{k=0}^{N-1}|D^{N-k}V||D^{k+1}F|+\sum_{k=0}^{N-1}|D^{k+1}F|\sum_{m=k+1}^{N}|D^{N-m}V|\Big{\}}$
$=$ $C\chi\sum_{k=1}^{N}|D^{k}F|\sum_{m=0}^{N-k+1}|D^{m}V|.$
Next, by (B.57a)–(B.57b), we have
$\nabla\textnormal{{div}}V$
$=$ $\hat{r}(\hat{r}\cdot\partial_{r}^{2}V)+\hat{\theta}\frac{\partial_{r}\partial_{\theta}(\hat{r}\cdot V)}{r}+\hat{\varphi}\frac{\partial_{r}\partial_{\varphi}(\hat{r}\cdot V)}{r\sin\theta}+2\hat{r}\Big{(}\frac{\partial_{r}(\hat{r}\cdot V)}{r}-\frac{\hat{r}\cdot V}{r^{2}}\Big{)}+\hat{\theta}\frac{2\partial_{\theta}(\hat{r}\cdot V)}{r^{2}}$
$+\hat{\varphi}\frac{2\partial_{\varphi}(\hat{r}\cdot V)}{r^{2}\sin\theta}+\hat{r}\Big{(}\frac{\partial_{r}\partial_{\theta}(\hat{\theta}\cdot V)}{r}-\frac{\partial_{\theta}(\hat{\theta}\cdot V)}{r^{2}}\Big{)}+\hat{\theta}\frac{\partial_{\theta}^{2}(\hat{\theta}\cdot V)}{r^{2}}+\hat{\varphi}\frac{\partial_{\varphi}\partial_{\theta}(\theta\cdot V)}{r^{2}\sin\theta}$
$+\hat{r}\Big{(}\frac{\partial_{r}(\hat{\theta}\cdot V)\cos\theta}{r\sin\theta}-\frac{(\hat{\theta}\cdot V)\cos\theta}{r^{2}\sin\theta}\Big{)}+\hat{\theta}\partial_{\theta}\Big{(}\frac{(\hat{\theta}\cdot V)\cos\theta}{r^{2}\sin\theta}\Big{)}+\hat{\varphi}\frac{\partial_{\varphi}(\hat{\theta}\cdot V)\cos\theta}{r^{2}\sin^{2}\theta}$
$+\hat{r}\Big{(}\frac{\partial_{r}\partial_{\varphi}(\hat{\varphi}\cdot V)}{r\sin\theta}-\frac{\partial_{\varphi}(\hat{\varphi}\cdot V)}{r^{2}\sin\theta}\Big{)}+\hat{\theta}\partial_{\theta}\Big{(}\frac{\partial_{\varphi}(\hat{\varphi}\cdot V)}{r^{2}\sin\theta}\Big{)}+\hat{\varphi}\frac{\partial_{\varphi}^{2}(\hat{\varphi}\cdot V)}{r^{2}\sin^{2}\theta}.$
Computing the commutator $[\partial^{k},\nabla\textnormal{{div}}]$ using the above expression, we obtain the last inequality of the proposition. $\square$
Proposition B.2.
Let $\partial\in\{\partial_{\theta},\partial_{\varphi}\}$ and $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$. Then for $k,l\in\mathbb{N}\cup\{0\}$, there exists a constant $C=C(k,l)>0$ such that, for all $V(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$,
$\int_{\Omega}\chi\big{|}\partial_{r}^{l}\partial^{k}(\hat{r}\cdot\Delta V-\partial_{r}\textnormal{{div}}V)\big{|}^{2}{\mathrm{d}}x\leq$ $C\int_{\Omega}\chi\sum_{i=1}^{l+1}\sum_{j=0}^{k+1}\big{|}D^{i}\partial^{j}V\big{|}^{2}{\mathrm{d}}x.$
Proof.
Using differential relations (B.56) and (B.57), we get
$\hat{r}\cdot\Delta V=$ $\hat{r}\cdot\partial_{r}^{2}V+\frac{2}{r}\hat{r}\cdot\partial_{r}V+\frac{\cot\theta}{r^{2}}\hat{r}\cdot\partial_{\theta}V+\frac{1}{r^{2}}\hat{r}\cdot\partial_{\theta}^{2}V+\frac{1}{r^{2}\sin^{2}\theta}\hat{r}\cdot\partial_{\varphi}^{2}V,$
$\partial_{r}\textnormal{{div}}V=$ $\hat{r}\cdot\partial_{r}^{2}V+\frac{\hat{\theta}\cdot\partial_{\theta}\partial_{r}V}{r}+\frac{\hat{\varphi}\cdot\partial_{\varphi}\partial_{r}V}{r\sin\theta}-\frac{\hat{\theta}\cdot\partial_{\theta}V}{r^{2}}-\frac{\hat{r}\cdot V}{r^{2}}-\frac{\cot\theta}{r^{2}}\hat{\theta}\cdot V.$
Taking the difference of the above two equations, we see that the term $\hat{r}\cdot\partial_{r}^{2}V$ cancels. Then, multiplying $\chi$ on both side of the resultant equality, integrating in $x\in\Omega$, and applying Proposition A.1 on terms of the form $\int_{\Omega}\frac{|V|^{2}}{r^{2}}{\mathrm{d}}x$, we obtain the desired inequality. $\square$